### v2

# Frank–Wolfe algorithm for DC optimization problem

Abstract
We propose two versions of the Frank–Wolfe (conditional gradient) algorithm with adaptive step sizes for DC optimization problems, where the objective is the difference of a differentiable convex function with Lipschitz continuous gradient and a convex (possibly non-differentiable) function. The second version builds on the first, using finite differences to approximate the gradient of the first component. Unlike previous approaches relying on curvature or Lipschitz constants, our stepsizes require no such parameters. We also introduce the class of piecewise-star-convex functions, which may be nonconvex and non–differentiable, and show that the first version achieves an $1/k$ convergence rate for minimizing such functions, while ensuring stationarity of all limit points. In the second version, the finite–difference scheme yields relative rather than absolute gradient errors, simplifying the analysis; we prove that all limit points are stationary and that the duality gap decreases at rate $1/\sqrt{k}$.
Keywords: Frank-Wolfe method; DC optimization problem; finite difference; weak-star-convex function.
AMS subject classification: 90C25, 90C60, 90C30, 65K05.
1 Introduction
The DC optimization method involves minimizing a function that can be represented as the difference between two convex functions. We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable. To the best of our knowledge, DC optimization can be traced back to pioneering works such as [42, 43], which introduced the first algorithms for this problem. Since then, the DC optimization has attracted the attention of the mathematical programming community (see for example [7, 11, 13, 32, 44]), not only for its own sake but also because it is an abstract model for several families of practical optimization problems. Although we are not concerned with practical issues at this time, we emphasize that practical applications emerge whenever the natural structure of the problem is modelled as a DC optimization problem, such as sparse generalized eigenvalue problems [41], sparse optimization problems [19], facility location and clustering problems [36].
The Frank-Wolfe algorithm has a long history, dating back to Frank and Wolfe’s work in the 1950s to minimize convex quadratic functions over compact polyhedral sets, see [15]. This method was generalized about ten years later to minimize convex differentiable functions with Lipschitz continuous gradients and compact constraint convex sets, see [33]. Since then, this method has attracted the attention of several researchers who work with continuous optimization, thus becoming also known as the conditional gradient method. One of the factors that explain the interest in this method is its simplicity and ease of implementation. In fact, each method iteration only requires access to a linear minimization oracle over a compact convex set. It is also worth noting that, due to the method’s simplicity, it allows for low storage costs and ready exploration of separability and sparsity, making its application in large-scale problems quite attractive. It is interesting to note that the popularity of this method has increased significantly in recent years as a result of the emergence of several applications in machine learning, see [25, 29, 30]. For all of these reasons, several variants of this method have arisen throughout the years and new properties of it have been discovered, resulting in a large literature on it, papers dealing with this method include [4, 8, 9, 17, 18, 21, 27, 31, 34].
In the present paper, we formulate two versions of Frank–Wolfe algorithm or conditional gradient method to solve DC optimization problem with an adaptive step size. The second version is based on the first and employs finite differences to approximate the gradient $\nabla g$. It is worth mentioning that the Frank-Wolfe algorithm has previously been formulated in the context of DC programming, see [26]. See also [46] which establishes theoretical connections between DC algorithms and convex-concave procedures with the Frank-Wolfe algorithm. In contrast to the previous study, which assume that the curvature/Lipschitz-type constant of f=ghf=g-h is bounded from above, the analysis of the Frank-Wolfe method done here just assumes that the gradient of the first component $g$ of $f$ is Lipschitz continuous. While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous in both formulations, we will not compute the step size using the Lipschitz constant of $\nabla g$, as in earlier formulations that employed the curvature/Lipschitz-type constant of the objective function $f$. The step size will be computed adaptively, based on an idea introduced in [3] (see also [5, 39]) that approximates the Lipschitz constant. It has been shown in previous works that Frank-Wolfe algorithm produces a stationary point with a convergence rate of only ${\cal O}(1/\sqrt{k})$ for non-convex objective functions, see [28] (see also [26]). We introduce the concept of weak-star-convex function notion, which generalizes the star-convex function concept proposed in [37] as well as a few other related concepts, such as [22, 45]. The weak-star-convex objective functions are also taken into account in the convergence analysis of the first method presented here. The convergence rate is $\mathcal{O}(1/k)$ for both the function values and the duality gap, even though piecewise-star-convexs are generally non-differentiable and non-convex. Consequently, the class of functions for which the Frank–Wolfe method achieves an $\mathcal{O}(1/k)$ convergence rate includes piecewise-star-convexs that are differences of convex functions, with the first component having a Lipschitz continuous gradient. This class complements the set of star–convex functions, which was recently shown in [35] to enjoy the same convergence rate. As mentioned previously, the second version of the Frank-Wolfe algorithm uses finite difference to approximate the gradient of the function $g$. Similar to the strategies adopted in [20], the finite difference utilized here to approximate the gradient is computed with the step-size updated adaptively utilizing two previous iterations. Finite difference is an old concept that has been used to approximate derivatives in optimization settings (for example, see [38, Section 8.1]) and has already appeared in the study of the Frank-Wolfe algorithm (for example, see [14, 16, 25, 40]). It is worth noting that all previous applications of finite difference in the Frank-Wolfe algorithm provided approximate gradients with absolute error, whereas the one used here has the advantage of providing us with a relative error and thus being quite simple to analyze. We show that all limit points of the generated sequence for the second version of the Frank-Wolfe algorithm are stationary points for the problem under consideration, and we establish that the rate of convergence is ${\cal O}(1/\sqrt{k})$ for the duality gap.
The remainder of the paper is organized as follows. Section 2 introduces the notation and collects auxiliary results. In Section 3 we define the concept of piecewise–star–convex, present its main properties, and provide illustrative examples. Section 4 states the optimization problem, the required assumptions, and related notation. Section 5 is devoted to the first version of the Frank–Wolfe algorithm, including its formulation, well–posedness, and analysis of the generated sequence. Section 6 presents the second version of the algorithm together with its formulation, well–posedness, and convergence analysis. Finally, Section 7 summarizes our conclusions.
2 Preliminaries
In this section, we recall some notations, definitions and basics results used throughout the paper. A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be convex if $\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)$, for all $x,y\in{\mathbb{R}^{n}}$ and $\lambda\in[0,1]$. And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex function see [23]. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz if, for all $x\in\mathbb{R}^{n}$, there exist a constant ${\cal C}_{x}>0$ and a neighborhood $U_{x}$ of $x$ such that $|f(x)-f(y)|\leq{\cal C}_{x}\|x-y\|$, for all $y\in U_{x}.$ It is well known that, if $f:\mathbb{R}^{n}\to\mathbb{R}$ is convex, then $f$ is locally Lipschitz. And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [10, p. 32]. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial_{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
$f^{\circ}(x;d)=\limsup_{\tiny{\begin{array}[]{c}u\rightarrow x\\ t\downarrow 0\end{array}}}\frac{f(u+td)-f(u)}{t}.$
For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [10, p. 27]. If $f$ is convex, then $\partial_{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [10, p. 36]. We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial_{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [10, p. 33].
Theorem 2.1 ([10, p. 27]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. Then, $\partial_{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial_{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$. Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial_{c}f(x)\}$.
Theorem 2.2 ([10, p. 38-39]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be given by f=ghf=g-h, where $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable. Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial_{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
The next result is a combination of Theorem 2.2 with [10, Corollary on p. 52].
Theorem 2.3.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function and ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a closed and convex set. If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and all $x\in{{\cal C}}$. As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $u\in\partial_{c}h({x^{*}})$ and all $x\in{{\cal C}}$.
Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$. A continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ has gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$ if there exists a Lipschitz constant $L>0$ such that $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$. Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [6, Proposition A.24], see also[12, Lemma 2.4.2].
Proposition 2.4.
Let $g:\mathbb{R}^{n}\to\mathbb{R}$ be a differentiable with gradient $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, $x\in{{\cal C}}$, $v\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$. If $x+\lambda v\in{{\cal C}}$, then $g(x+\lambda v)\leq g(x)+\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}v\lambda+\frac{L}{2}\|v\|^{2}\lambda^{2}.$
Proposition 2.5.
The function $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex if and only if $h(y)\geq h(x)+\langle u,y-x\rangle$, for all $x,y\in\mathbb{R}^{n}$ and all $u\in\partial h(x)$.
Proposition 2.6 ([23, Proposition 6.2.1]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. Let $(x^{k})_{k\in{\mathbb{N}}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$. If $\lim_{k\to+\infty}x^{k}={\bar{x}}$ and $\lim_{k\to+\infty}u^{k}={\bar{u}}$, then ${\bar{u}}\in\partial h({\bar{x}})$.
Proposition 2.7 ([23, Proposition 6.2.2]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. The mapping $\partial h$ is locally bounded, i.e. the image $\partial h(B)$ of a bounded set $B\subset\mathbb{R}^{n}$ is a bounded set in $\mathbb{R}^{n}$.
In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [2, Lemma 13.13, Ch. 13, p. 387].
Lemma 2.8.
Let $(a_{k})_{k\in{\mathbb{N}}}$ and $(b_{k})_{k\in{\mathbb{N}}}$ be nonnegative sequences of real numbers satisfying
$a_{k+1}\leq a_{k}-b_{k}\beta_{k}+\frac{A}{2}\beta_{k}^{2},\qquad k=0,1,2,\ldots,$
where $\beta_{k}=2/(k+2)$ and $A$ is a positive number. Suppose that $a_{k}\leq b_{k}$, for all $k$. Then
$a_{k}\leq\frac{2A}{k}$, for all $k=1,2,\ldots.$
$\min_{\ell\in\{\lfloor\frac{k}{2}\rfloor+2,\ldots,k\}}b_{\ell}\leq\frac{8A}{k-2}$, for all $k=3,4,\ldots,$ where, $\lfloor k/2\rfloor=\max\left\{n\in\mathbb{N}:~n\leq k/2\right\}.$
We conclude this section by recalling the definitions of some norms that will be used later. For $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$ and $p\in[1,+\infty)$, the pp–norm is $\|x\|_{p}:=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{\!1/p}$. In particular, the $\ell_{1}$–norm is $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, and the $\ell_{\infty}$–norm is $\|x\|_{\infty}:=\max_{1\leq i\leq n}|x_{i}|$.
3 Piecewise-star-convex functions
In this section, we introduce the concept of piecewise-star-convexity, a relaxation of the classical star–convexity notion from [37]. Many DC functions that are nonconvex nevertheless satisfy this weaker property. As will be shown in a later section, by identifying and exploiting piecewise–star–convexity we can broaden the class of problems for which the classical Frank–Wolfe algorithm attains an ${\cal O}(1/k)$ convergence rate. We begin by recalling the definition of star–convexity from [37] and then present its relaxed counterpart, the piecewise-star-convex notion, which will be central to our analysis.
Definition 3.1.
Let ${{\cal C}}\subset{\mathbb{R}}^{n}$ be a convex set. A function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ is star–convex on $\cal C$}} if its set of global minimizers $X^{*}:=\arg\min_{{\cal C}}f$ is nonempty and, for any $x^{*}\in X^{*}$,
$f(\lambda x^{*}+(1-\lambda)x)\leq\lambda f(x^{*})+(1-\lambda)f(x),\qquad\forall\,x\in{{\cal C}},\ \forall\,\lambda\in[0,1].$ (1)
Every convex function with a nonempty set of global minimizers is star–convex, but the converse does not hold. We present two nonconvex star–convex functions from [37] (see also [35] for additional examples) to highlight that they are DC functions, each expressible as the difference of two convex functions.
Example 3.2.
The star–convex function $\phi(t)=|t|(1-e^{-|t|})$ is not convex. In addition, $\phi$ is a difference of two convex functions. Indeed, considering the following two convex functions $\varphi(t)=|t|(1-e^{-|t|})+e^{|t|}$ and $\psi(t)=e^{|t|}$, we have $\phi=\varphi-\psi$. Similarly, consider the star–convex function $f(s,t)=s^{2}t^{2}+s^{2}+t^{2}$. The function $f$ is not convex, but is a difference of two convex functions. Indeed, letting the following convex functions $g(s,t)=(s^{2}+t^{2})^{2}+s^{2}+t^{2}$ and $h(s,t)=s^{2}t^{2}+s^{4}+t^{4}$, we have f=ghf=g-h.
The requirement in (1) is formulated with respect to a fixed minimizer $x^{*}$ and must hold for all $x\in{\cal C}$. A milder, yet still powerful, relaxation is to allow the reference minimizer to depend on $x$. In many DC models, the reference minimizer that makes (1) valid for a given $x$ may vary with $x$, which is precisely what piecewise-star-convexity captures. Even this global condition can be too restrictive when $f$ exhibits multiple “basins” of favorable geometry. A natural remedy is to partition the domain into regions within which a local star–type inequality holds with respect to a (local) minimizer:
Definition 3.3.
Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets with pairwise disjoint interiors and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$} if, for every $i\in I$,
  • (i)
    the set of global minimizers of $f$ on ${\cal C}_{i}$, denoted $X_{i}^{*}$, is nonempty; and
  • (ii)
    for each $x\in{\cal C}_{i}$ there exists $x_{x}^{*}\in X_{i}^{*}$ such that
    $f\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x),\qquad\forall\,\lambda\in[0,1].$ (2)
Remark 1.
When the partition is trivial (${\cal P}=\{{\cal C}\}$) and $X_{1}^{*}=X^{*}$, Definition 3.3 reduces to Definition 3.1. In practice, piecewise-star-convexity often arises when $f$ is convex on each region ${\cal C}_{i}$ but not across different regions. In this setting, the minimizers $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ are global minimizers within their corresponding cells ${\cal C}_{i}$, although they may be only local (and possibly non–global) minimizers of $f$ on the whole domain $\cal C$}. This structure ensures that star–type inequalities hold on each cell (with respect to its minimizers $X_{i}^{*}$), even if such inequalities fail when crossing cell boundaries.
We emphasize that in Definition 3.3, the collection of subsets in item (ii) need not be finite: the piecewise-star-convexity property may still hold when $\cal C$} is covered by infinitely many convex regions. The following example illustrates that finiteness cannot, in general, be assumed.
Example 3.4.
Consider the set ${\cal C}=[0,1]$ and the infinite set $S:=\{0\}\cup\{1/k:k\in\mathbb{N}\}\subset[0,1],$ which is compact, infinite, and has an accumulation point at 0. Define $f:[0,1]\to\mathbb{R}$ by
$f(x)=\min_{c\in S}|x-c|^{2},\qquad x\in[0,1],$
see the Figure 1. The set of global minimizers is $X^{*}=\operatorname*{argmin}_{x\in[0,1]}f(x)=S$ with $\min f=0$. For each $x\in[0,1]$, choose $x_{x}^{*}\in S$ attaining the minimum in the definition of $f$, and let
$V_{x_{x}^{*}}:=\{\,y\in[0,1]:|y-x_{x}^{*}|\leq|y-c|\ \forall c\in S\,\}.$
These sets form a countable partition $0,1]=\bigcup_{k=0}^{\infty}V_{k}$, with
$V_{1}=\Bigl{[}\tfrac{3}{4},\,1\Bigr{]},\quad V_{k}=\Bigl{[}\tfrac{1/k+1/(k+1)}{2},\ \tfrac{1/k+1/(k-1)}{2}\Bigr{]},\ k\geq 2,\quad V_{0}=\{0\}.$
By construction, $x\in V_{x_{x}^{*}}$ and the whole segment $x_{x}^{*},x]$ lies in $V_{x_{x}^{*}}$. Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y\in[x_{x}^{*},x]$. Hence, for any $\lambda\in[0,1]$,
$f\big{(}\lambda x_{x}^{*}+(1-\lambda)x\big{)}=\big{|}\lambda x_{x}^{*}+(1-\lambda)x-x_{x}^{*}\big{|}^{2}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x).$
Since $f(x_{x}^{*})=0$, the piecewise-star-convexity inequality holds. However, covering $\cal C$} by convex sets satisfying (i)–(ii) requires all sets $V_{k}$ above, so any such cover must be infinite.
[…]
Figure 1: Plot of the function $f(x)=\min_{c\in S}(x-c)^{2}$ for $S=\{0,\tfrac{1}{8},\tfrac{1}{7},\tfrac{1}{6},\tfrac{1}{5},\tfrac{1}{4},\tfrac{1}{3},\tfrac{1}{2},1\}$. The black segments indicate the active arc of the lower envelope, the dashed lines mark the boundaries of $V(x_{x}^{*})$, and the solid points on the $x$-axis represent the centers $c\in S$.
We next present a prototypical example of a function that is piecewise-star-convex. This model arises naturally in sparse estimation, robust quadratics, and regularized control. The $\ell_{1}$ penalty acts as a sign–dependent linear perturbation of a positive–definite quadratic. By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the function reduces on each region to a strictly convex quadratic with a unique minimizer. These cell minimizers provide the reference points in Definition 3.3, thereby establishing piecewise-star-convexity even though, globally, such minimizers may be only local.
For use in the examples below and to simplify notation, define
$\{-1,1\}^{n}=\{s:=(s_{1},\ldots,s_{n})\in\mathbb{R}^{n}:~s_{i}\in\{-1,1\},~\forall i\in\{1,\ldots,n\}\}.$ (3)
Example 3.5.
Let $Q\in\mathbb{R}^{n\times n}$ be positive definite and let $\beta>0$. Consider $f:{\mathbb{R}}^{n}\to\mathbb{R}$ given by
$f(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$ (4)
For each $s\in\{-1,1\}^{n}$, define the closed orthant
$\Omega_{s}=\bigl{\{}\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0,\,\text{for }i=1,\ldots,n\bigr{\}},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
On each $\Omega_{s}$, the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$, for $x\in\Omega_{s}$. Hence, the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$
with constant Hessian $Q$. Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\mathbb{R}^{n}$ with modulus $\lambda_{\min}$, the minimizer $x_{s}^{*}$ is unique. Considering that $f$ coincides with the strictly convex quadratic $f_{s}$ on each orthant $\Omega_{s}$ and $x_{s}^{*}$ is the (unique) minimizer of $f$ on $\Omega_{s}$, the convexity inequality
$f\bigl{(}\lambda x_{s}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{s}^{*})+(1-\lambda)f(x),\qquad\forall\,x\in\Omega_{s},\ \forall\,\lambda\in[0,1],$
holds. Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$.
Next we preste a constrained DC optimization problem motivated by contrastive learning formulations.
Example 3.6.
Given two disjoint sets of reference points, a set of positive examples $\mathcal{P}=\{\bar{p}_{1},\dots,\bar{p}_{m}\}\subset\mathbb{R}^{n}$ and a set of negative examples $\mathcal{N}=\{\bar{q}_{1},\dots,\bar{q}_{r}\}\subset\mathbb{R}^{n}$, the objective is to identify a point $x$ that lies close to the positive references while remaining far from the negative ones. Here all distances are measured with respect to the standard Euclidean metric. To remain consistent with our general framework, we impose that the search is carried out over a closed and convex feasible set $\mathcal{C}\subset\mathbb{R}^{n}$. Formally, we study the problem
$\min_{x\in\mathcal{C}}f(x),$ (5)
where the Euclidean contrastive objective is defined by
$f(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}-\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},\qquad x\in\mathbb{R}^{n},$ (6)
with $\lambda_{i}^{+}>0$ and $\lambda_{j}^{-}>0$ fixed weights. The squared $\ell_{2}$ terms pull $x$ toward the positive examples, while the negative $\ell_{1}$ term acts as a contrastive regularizer, promoting separation from the negative examples in a coordinate-dependent way. The objective (6) admits a natural DC decomposition of the form
$g(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2},\qquad h(x):=\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},$ (7)
where $g$ is smooth and strongly convex, while $h$ is convex but nondifferentiable. Hence similarly to Example 3.5 we can shoe that f=ghf=g-h belongs to the class of piecewise-star-convex DC models. In fact, let $m\geq 1$ and $\lambda_{i}^{+}>0$ for $i=1,\ldots,m$, and define $\Lambda^{+}:=\sum_{i=1}^{m}\lambda_{i}^{+}>0$. Consider $f:\mathbb{R}^{n}\to\mathbb{R}$ given in (6). Then $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the partition ${\cal P}=\{{\cal C}_{S}\}_{S}$ induced by the hyperplane arrangement $\{x_{k}=\bar{q}_{j,k}:~j=1,\ldots,r,\ k=1,\ldots,n\}$, where for $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$,
${\cal C}_{S}:=\bigl{\{}\,x\in\mathbb{R}^{n}:\ s_{j,k}\,(x_{k}-\bar{q}_{j,k})\geq 0,\ \forall j=1,\ldots,r,\ \forall k=1,\ldots,n\,\bigr{\}}.$
Fix $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$ and consider $x\in{\cal C}_{S}$. For every $j$ and $k$ we have $|x_{k}-\bar{q}_{j,k}|=s_{j,k}(x_{k}-\bar{q}_{j,k})$ on ${\cal C}_{S}$, hence
$\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1}\;=\;\sum_{j=1}^{r}\lambda_{j}^{-}\sum_{k=1}^{n}s_{j,k}(x_{k}-\bar{q}_{j,k})\;=\;v_{S}^{\top}x-c_{S},$
where $v_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}\in\mathbb{R}^{n}$ and $c_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}^{\top}\bar{q}_{j}$ is a constant on ${\cal C}_{S}$. Moreover,
$\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}\;=\;\Lambda^{+}\|x\|_{2}^{2}-2\,b^{\top}x\;+\;d,$
with $b:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\bar{p}_{i}$ and $d:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|\bar{p}_{i}\|_{2}^{2}$. Therefore, on ${\cal C}_{S}$ the function $f$ reduces to the strongly convex quadratic
$f_{S}(x)\;=\;\Lambda^{+}\|x\|_{2}^{2}-(2b+v_{S})^{\top}x\;+\;(d-c_{S}),$
whose Hessian is $2\Lambda^{+}I$ and is positive definite. Since ${\cal C}_{S}$ is closed and convex and $f_{S}$ is strongly convex, $f_{S}$ admits a unique minimizer $x_{S}^{\ast}\in{\cal C}_{S}$. By convexity of $f_{S}$ on the convex set ${\cal C}_{S}$, for all $x\in{\cal C}_{S}$ and all $\lambda\in[0,1]$,
$f\bigl{(}\lambda x_{S}^{\ast}+(1-\lambda)x\bigr{)}\;\leq\;\lambda f(x_{S}^{\ast})+(1-\lambda)f(x).$
This is precisely the star–type inequality required in Definition 3.3 with anchor $x_{S}^{\ast}$ on the cell ${\cal C}_{S}$. Since the interiors of the cells $\{{\cal C}_{S}\}$ are pairwise disjoint and $\bigcup_{S}{\cal C}_{S}=\mathbb{R}^{n}$, the claim follows.
The structure f=ghf=g-h reflects the central principle of contrastive learning: the quadratic component promotes similarity with positive references, while the $\ell_{1}$ component enforces dissimilarity from negative ones. For comprehensive discussions of contrastive learning in the Euclidean setting, see [24]. In the next proposition, we establish that, under a suitable dominance condition on the positive weights, the function $f$ in (6) is coercive and therefore attains at least one global minimizer of (5).
The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization (see [Libertietall2014] and [1] for a comprehensive survey on this subject). This construction is also connected to Example 3.4.
Example 3.7.
Let $Q\in\mathbb{R}^{n\times n}$ be positive definite, and let $b_{1},\dots,b_{m}\in\mathbb{R}^{n}$ and $c_{1},\dots,c_{m}\in\mathbb{R}$. Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
$f(x):=\min_{1\leq i\leq m}\varphi_{i}(x),\qquad\qquad\varphi_{i}(x):=\tfrac{1}{2}\,x^{\top}Qx+b_{i}^{\top}x+c_{i},\qquad i=1,\dots,m.$ (8)
Consider the partition ${\cal P}:=\{{\cal C}_{i}:~i=1,\ldots,m\}$ given by
${\cal C}_{i}=\bigl{\{}x\in\mathbb{R}^{n}:\ \varphi_{i}(x)\leq\varphi_{j}(x)\ \ \forall j=1,\dots,m\bigr{\}}=\bigcap_{j=1}^{m}\bigl{\{}x\in\mathbb{R}^{n}:\ (b_{i}-b_{j})^{\top}x\leq c_{j}-c_{i}\bigr{\}},$
since $\varphi_{i}(x)-\varphi_{j}(x)=(b_{i}-b_{j})^{\top}x+(c_{i}-c_{j})$. Hence each ${\cal C}_{i}$ is a closed convex polyhedron. By construction, we have
$f(x)=\varphi_{i}(x)\qquad\forall\,x\in{\cal C}_{i}.$
Because $Q$ is positive definite, each $\varphi_{i}$ is strongly convex; therefore $f$ coincides with a strongly convex quadratic on each cell ${\cal C}_{i}$. In particular, $f$ is convex on each ${\cal C}_{i}$, hence star–convex there, and thus piecewise star–convex with respect to the polyhedral partition $\{{\cal C}_{i}\}_{i=1}^{m}$.
Next, observe that $f$ admits a DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
$\min_{1\leq i\leq m}\Bigl{\{}\tfrac{1}{2}\,x^{\top}Qx+b_{i}^{\top}x+c_{i}\Bigr{\}}=\tfrac{1}{2}\,x^{\top}Qx+\min_{1\leq i\leq m}\bigl{\{}b_{i}^{\top}x+c_{i}\bigr{\}}=\tfrac{1}{2}\,x^{\top}Qx-\max_{1\leq i\leq m}\bigl{\{}-b_{i}^{\top}x-c_{i}\bigr{\}}.$
Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx)$ and $h(x):=\max_{1\leq i\leq m}\bigl{(}-b_{i}^{\top}x-c_{i}\bigr{)}$. Clearly $g$ is $C^{\infty}$ and convex, while $h$ is the pointwise maximum of affine functions, hence convex, polyhedral, and Lipschitz.
A particularly relevant special case of (8) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $f(x)=\min_{1\leq i\leq m}\,\|x-a_{i}\|^{2},\qquad x\in\mathbb{R}^{n}.$
Next we show that a basic robustness property of piecewise–star–convexity is that the star–type inequality is preserved by adding a convex term.
Proposition 3.8.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}. Then $f+g$ is piecewise–star–convex on $\cal C$} with respect to $\cal P$}.
Proof.
Fix $i\in I$ and $x\in{\cal C}_{i}$. By Definition 3.3, there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ such that
$f\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x),\quad\forall\,\lambda\in[0,1].$
Since $g$ is convex on ${\cal C}_{i}$,
$g\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x),\quad\forall\,\lambda\in[0,1].$
Summing the inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$. As the argument holds for every $i\in I$ and $x\in{\cal C}_{i}$, the claim follows. ∎
Remark 2.
If your definition of piecewise–star–convexity explicitly requires the reference points to be minimizers of the function under consideration on each cell, then the statement above remains valid provided $(f+g)$ attains a (local) minimum on each ${\cal C}_{i}$ (e.g., when each ${\cal C}_{i}$ is compact or $(f+g)$ is coercive there). In that case one may choose the reeference points among $\arg\min_{{\cal C}_{i}}(f+g)$.
The next example combines Example 3.5 with Proposition 3.8 to show that a widely used optimization objective is piecewise–star–convex.
Example 3.9.
Let $A\in\mathbb{R}^{n\times n}$ be a full rank column, equivalently, $Q:=A^{\top}A$ a positive definite matrix, let $b\in\mathbb{R}^{n}$, and let $\beta>0$. Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{\,2}-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$
Expanding the quadratic term gives the equivalent representation
$f(x)\;=\;\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}-c^{\top}x\;+\;\tfrac{1}{2}\|b\|_{2}^{2},\qquad Q:=A^{\top}A,\qquad c:=A^{\top}b.$ (9)
Equivalently, write $f=h+g$ with $h(x):=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}$ and $g(x):=-\,c^{\top}x+\tfrac{1}{2}\|b\|_{2}^{2}$. By Example 3.5, $h$ is piecewise–star–convex on the orthant partition when $Q\succ 0$; since $g$ is convex (affine), Proposition 3.8 implies that $f=h+g$ is also piecewise–star–convex on the same partition.
Proposition 3.10.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is closed and convex, interiors are pairwise disjoint, and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$. Let $A\in\mathbb{R}^{n\times n}$ be invertible and $b\in\mathbb{R}^{n}$. Define the affine map
$\Psi(z):=Az+b,\qquad z\in\mathbb{R}^{n},$
the transformed domain and cells $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
$\tilde{f}(z):=f(Az+b),\qquad z\in\tilde{\mathcal{C}}$.
Then $\tilde{f}$ is piecewise–star–convex on $\tilde{\mathcal{C}}$ with respect to the partition $\tilde{\mathcal{P}}=\{\tilde{\mathcal{C}}_{i}\}_{i\in I}$.
Proof.
(i) Since $A$ is invertible, $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$. If $z\in\tilde{\mathcal{C}}$ then $x:=\Psi(z)=Az+b\in\mathcal{C}$, so $\tilde{f}(z)=f(x)$ is well defined. Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is closed and convex; interiors remain pairwise disjoint and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$. Thus $\tilde{\mathcal{P}}$ is a partition of $\tilde{\mathcal{C}}$.
(ii) For each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$. We claim
$\tilde{X}_{i}^{*}=A^{-1}(X_{i}^{*}-b).$ (10)
Indeed, for $z\in\tilde{\mathcal{C}}_{i}$ write $x=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Then $\tilde{f}(z)=f(x)$ and hence
$\min_{z\in\tilde{\mathcal{C}}_{i}}\tilde{f}(z)=\min_{x\in\mathcal{C}_{i}}f(x),\qquad z\in\tilde{X}_{i}^{*}\ \Longleftrightarrow\ x=\Psi(z)\in X_{i}^{*}.$
Equivalently, $z=A^{-1}(x-b)$ with $x\in X_{i}^{*}$, proving (10).
(iii) Fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Because $f$ is piecewise–star–convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
$f\big{(}(1-\lambda)x+\lambda x_{x}^{*}\big{)}\leq(1-\lambda)f(x)+\lambda f(x_{x}^{*}).$ (11)
Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (10), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$,
$\Psi\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}=(1-\lambda)\Psi(z)+\lambda\Psi(z_{z}^{*})=(1-\lambda)x+\lambda x_{x}^{*}.$
Therefore,
$\tilde{f}\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}$ $=f\!\left(\Psi\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}\right)=f\big{(}(1-\lambda)x+\lambda x_{x}^{*}\big{)}$
$\leq(1-\lambda)f(x)+\lambda f(x_{x}^{*})=(1-\lambda)\tilde{f}(z)+\lambda\tilde{f}(z_{z}^{*}),$
where the inequality uses (11) and the identities $f(x)=\tilde{f}(z)$, $f(x_{x}^{*})=\tilde{f}(z_{z}^{*})$. Hence, for every $z\in\tilde{\mathcal{C}}_{i}$ there exists $z_{z}^{*}\in\tilde{X}_{i}^{*}$ such that the star–type inequality holds on $\tilde{\mathcal{C}}_{i}$. Since $i$ was arbitrary, $\tilde{f}$ is piecewise–star–convex on $\tilde{\mathcal{C}}$ with respect to $\tilde{\mathcal{P}}$. ∎
The next result establishes a first-order inequality that holds for any piecewise-star-convex. This property will play a key role in the convergence analysis of our algorithm.
Proposition 3.11.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz on a convex set $\cal C$}. Assume $f$ is piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$ of Definition 3.3. Fix $i\in I$ and $x\in{\cal C}_{i}$. Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ (a minimizer on the cell ${\cal C}_{i}$; hence, a local minimizer for $f$ on $\cal C$}) such that
$f(x_{x}^{*})-f(x)\ \geq\ f^{\circ}(x;\,x_{x}^{*}-x)\ \geq\ v^{\top}(x_{x}^{*}-x)\qquad\forall\,v\in\partial_{c}f(x),$ (12)
where $f^{\circ}$ denotes Clarke’s generalized directional derivative and $\partial_{c}f(x)$ the Clarke subdifferential at $x$.
Proof.
By piecewise–star–convexity on ${\cal C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that
$f\bigl{(}x+\lambda(x_{x}^{*}-x)\bigr{)}\ \leq\ (1-\lambda)f(x)+\lambda f(x_{x}^{*}),\qquad\forall\,\lambda\in[0,1].$
Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus (e.g., Theorem 2.1), $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (12). ∎
The next result shows that, under piecewise-star-convexity, every Clarke stationary point of the constrained problem is in fact a minimizer of $f$ on the cell to which it belongs.
Corollary 3.12.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$. Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary point of the constrained problem $\min_{x\in{\cal C}}f(x),$ that is,
$v^{\top}(x-\bar{x})\ \geq\ 0\qquad\forall\,v\in\partial_{c}f(\bar{x}),\ \forall\,x\in{\cal C}.$ (13)
If, in addition, $\bar{x}\in{\cal C}_{i_{0}}$ for some $i_{0}\in I$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{0}}$, namely, $\bar{x}\in X_{i_{0}}^{*}:=\arg\min_{x\in{\cal C}_{i_{0}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{0}}}f(x).$
Proof.
Apply Proposition 3.11 with $x=\bar{x}$ and $i=i_{0}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ with
$f(x_{\bar{x}}^{*})-f(\bar{x})\ \geq\ v^{\top}(x_{\bar{x}}^{*}-\bar{x})\qquad\forall\,v\in\partial_{c}f(\bar{x}).$
By stationarity (13) (taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{0}}\subset{\cal C}$), the right-hand side is $\geq 0$, hence $f(x_{\bar{x}}^{*})\geq f(\bar{x})$. Since $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ minimizes $f$ on ${\cal C}_{i_{0}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$. Therefore $f(x_{\bar{x}}^{*})=f(\bar{x})$ and $\bar{x}\in X_{i_{0}}^{*}$. ∎
Remark 3.
In the piecewise setting, $X_{i}^{*}$ consists of the global minimizers of $f$ restricted to the cell ${\cal C}_{i}$, which are in general only local minimizers on the whole set $\cal C$}. Corollary 3.12 refines the global statement for piecewise-star-convex functions by showing that Clarke stationarity implies cellwise optimality. Moreover, if the cell ${\cal C}_{i_{0}}$ containing $\bar{x}$ attains the global minimum value, i.e., $\min_{{\cal C}_{i_{0}}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}. This highlights the flexibility of piecewise-star-convexity: by allowing the domain to be partitioned into regions where the function behaves “locally convex,” one captures a broader class of DC objectives than with classical star–convexity, while still retaining descent inequalities sufficient to establish convergence of Frank–Wolfe variants.
4 The DC optimization problem
We are interested in solving the following constrained DC optimization problem
$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (14)
where ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact and convex set, $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function and $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex function possibly non-differentiable, with domain on $\mathbb{R}^{n}$. Throughout the paper we assume that the gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, i.e., there exists a Lipschitz constant $L>0$ such that
  • (A)
    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
Since we are assuming that ${\cal C}\subset{\mathbb{R}}^{n}$ is a compact set, its diameter is a finite number defined by
$\operatorname{diam}({\cal C}):=\max\left\{\|x-y\|:~x,y\in{\cal C}\right\}.$ (15)
Since ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact, the study of problem (14) is bounded from below. Then, the optimum value of the problem (14) satisfies $\infty<f^{*}:=\inf_{x\in{{\cal C}}}f(x)$ and optimal set $\cal C$$^{*}$ is non-empty. According to Theorem 2.3, the first-order optimality condition for problem (14) is stated as
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(x-{\bar{x}})\geq 0,\qquad\forall{\bar{u}}\in\partial g({\bar{x}}),\quad\forall x\in{{\cal C}}.$ (16)
In general, the condition (16) is necessary but not sufficient for optimality. A point ${\bar{x}}\in{{\cal C}}$ satisfying condition (16) is called a stationary point to problem (14). Consequently, all $x^{*}\in{{\cal C}}^{*}$ satisfies (16). We finish this section by presenting a variant of a classical example of DC optimization problem that fits the aforementioned requirements.
Example 4.1.
Let ${\cal C}_{i}\subset{\mathbb{R}}^{n}$ be an arbitrary set, for all $i=1,\ldots,m$. The square distance with respect to the set ${\cal C}_{i}$, denoted by $d^{2}_{{\cal C}_{i}}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is defined
$d^{2}_{{\cal C}_{i}}(x):=\inf_{y\in{\cal C}_{i}}\|x-y\|^{2}.$ (17)
In general, the function $d^{2}_{{\cal C}_{i}}$ is not convex. However, due to $\|x-y\|^{2}=\|x\|^{2}-2x^{T}y+\|y\|^{2}$, the distance function $d^{2}_{{\cal C}_{i}}$ becomes to
$d^{2}_{{\cal C}_{i}}(x):=\|x\|^{2}-\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right),$ (18)
which is a difference of two convex functions, see [23, Example 2.1.4]. Indeed, $d^{2}_{{\cal C}_{i}}=g-h$, where $g(x)=\|x\|^{2}$ is a convex quadratic function and $h(x)=\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)$ is a convex function given by the supremum of affine functions $x\mapsto\ell_{y}(x):=2x^{T}y-\|y\|^{2}$, for $y\in{\cal C}_{i}$. Let us state the constrained generalized Fermat-Weber location problem with square distances. For that, take $\omega_{i}\geq 0$, $i=1,\ldots,m$, such that $\sum_{i=1}^{m}\omega_{i}=1$ and note that
$\sum_{i=1}^{m}\omega_{i}d^{2}_{{\cal C}_{i}}(x)=\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right).$
Thus, for a given constrained set compact and convex ${\cal C}\subset{\mathbb{R}}^{n}$, the DC version of the constrained generalized Fermat-Weber location problem is stated as follows
$\begin{array}[]{c}\min_{x\in{{\cal C}}}\left(\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)\right).\end{array}$ (19)
Finally, observe that the objective function in problem (19) satisfies all the assumptions of problem 14. Moreover, in certain cases the objective function of the generalized Fermat–Weber problem (19) is weak–star–convex. For example, if $m=n$ and ${\cal C}_{i}=\{e_{i},-e_{i}\}$, where $e_{i}$ denotes the $i$-th canonical vector of $\mathbb{R}^{n}$, then the objective becomes $f(x)=\|x\|^{2}-2\sum_{i=1}^{n}\omega_{i}|x_{i}|+1,$ which, by the same reasoning as in Example 3.5, can be shown to be weak–star–convex.
5 Classical Frank–Wolfe algorithm with an adaptive stepsize
In this section, we present the classic Frank–Wolfe algorithm with an adaptive stepsize to solve problem (14) and provide its convergence analysis. Although we assume condition (A) for $g$, the Lipschitz constant will not be used explicitly to determine the stepsize. Instead, the stepsize is computed adaptively using the scheme introduced in [3, 5], which estimates the Lipschitz constant on the fly. Our convergence analysis shows that, for piecewise-star-convexfunctions, possibly non–differentiable and non–convex, the method achieves an ${\cal O}(1/k)$ convergence rate for both the function values and the duality gap. It is worth noting that a variant of the classic Frank–Wolfe algorithm for solving problem (14) was proposed in [26], where the stepsize depends on a curvature/Lipschitz–type constant of the gradient of $f$. For non–convex objective functions, the convergence analysis in [28] shows that such a variant attains only an ${\cal O}(1/\sqrt{k})$ rate in terms of reaching a stationary point.
Throughout this section, we assume access to a linear optimization oracle (LO oracle) that can minimize linear functions over the set $\cal C$}. The algorithm is stated below.
Algorithm 1.
Frank–WolfeC,f:=g-h algorithm
Step 0.
Select $x^{0}\in{{\cal C}}$ and $L_{0}>0$. Set $k=0$.
Step 1.
Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{0}\}$.
Step 2.
Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}(x_{k})$ as follows
$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}(x_{k}):=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (20)
Step 3.
If ${\omega}(x_{k})=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}(x_{k})|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (21)
Step 4.
If
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (22)
then set $j_{k}=j$ and go to Step 5. Otherwise, set $j=j+1$ and go to Step 3.
Step 5.
Set $\lambda_{k}:=\lambda_{j_{k}}$ and define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
$x^{k+1}:=x^{k}+\lambda_{k}(p^{k}-x^{k}),\qquad\quad L_{k+1}:=2^{j_{k}-1}L_{k}.$ (23)
Set $k\leftarrow k+1$, and go to Step 1.
Let us describe the main features of Algorithm 1. First of all, note that due to $x^{0}\in{C}$, $p^{k}\in{C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and C{C} be convex set, it follows from inductive arguments and first equality in (23) that $(x^{k})_{k\in{\mathbb{N}}}\subset{C}$. Note that (20) implies that
${\omega}(x_{k})\leq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall p\in{C}.$ (24)
Consequently, if ${\omega}(x_{k})=0$, then $(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0$, for all $p\in{C}$, i.e., $x^{k}$ is a stationary point to problem (14). Note that letting $p=x^{k}$ in (24) we conclude that
${\omega}(x_{k})\leq 0,\qquad k=0,1,\ldots.$ (25)
In order to simplify the notations, from now on we will use the following two notations:
${\omega}_{k}:={\omega}(x_{k}).$ (26)
Since $\omega_{k}=0$ implies that $x^{k}$ is a stationary point to problem (14). Thus, in view of (25) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$. The next result shows that the backtracking line–search in Algorithm 1 is well defined, identifies the smallest $j_{k}$ in Step 4, and establishes a descent inequality for the sequence $(x^{k})_{k\in{\mathbb{N}}}$.
Proposition 5.1.
For all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, the inequality (22) holds. Consequently, the number $j_{k}$ in Step 4 is well defined. Furthermore, $j_{k}$ is the smallest non-negative integer satisfying the following two conditions
$2^{j_{k}}L_{k}\geq 2L_{0},$ (27)
$f(x^{k}+\lambda_{k}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}_{k}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}.$ (28)
In addition, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. And the following inequality holds
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k},\qquad\qquad k=0,1,\ldots.$ (29)
Proof.
First we will prove that (22) holds for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$. It follows from Proposition 2.4 that
$g(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}.$ (30)
On the other hand, considering that $u^{k}\in\partial h(x^{k})$, it follows from Proposition (2.5) that
$h(x^{k}+\lambda_{j}(p^{k}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{k}-x^{k}\rangle.$ (31)
Thus, combining (30) with (31) and taking into account the equality in (20) and (25), some algebraic manipulations shows that
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega_{k}}|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L-{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}.$
Hence, it follows from the last inequality that (22) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proofs the first statement. Consequently, $j_{k}$ is well defined and is the smallest non-negative integer satisfying (27) and (28). Therefore, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. We proceed with the proof of (29), by showing the following inequality
$|{\omega_{k}}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}\leq-\frac{1}{2}|{\omega_{k}}|\lambda_{k}.$ (32)
For that we separately consider two cases: $\lambda_{k}=|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$ and ${j_{k}}=1$. In the former, we have $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}=|{\omega_{k}}|/{\lambda_{k}}$,which substituting into the left hand side of (32) yields an equality. If now $\lambda_{k}=1$, then it follows from (21) that $\lambda_{k}=1\leq|{\omega_{k}}|/({2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2})$ or equivalently ${2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|$ and $\lambda_{k}=1$, which again substituting into the left hand side of (32) yields the inequality. Hence, (32) holds. Therefore, (29) follows by combining (28) with (32), which concludes the proof. ∎
For simplifying the notations we define the following constants
$\alpha:=2(L+L_{0})\operatorname{diam}({{\cal C}})^{2}>0.$ (33)
Lemma 5.2.
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{0}\leq L_{k}\leq L+L_{0},\quad\qquad k=0,1,\ldots.$ (34)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{\alpha}\right\},\quad\qquad k=0,1,\ldots.$ (35)
Proof.
We proceed by induction to prove (34). Since the Lipschitz constant $L>0$, (34) holds for $k=0$. Now, assume that (34) holds for some $k>0$. We separately consider two cases: $j_{k}=0$ and $j_{k}\geq 1$. First assume that $j_{k}=0$. In this case, Step 1 together with second equality in (23) imply that $L_{k+1}=2^{-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}=0$. Since we are under the assumption that (34) holds, we have $L_{k+1}=2^{-1}L_{k}<L_{k}\leq L+L_{0}$. Thus, we conclude that two inequalities in (34) holds for $k+1$ and $j_{k}=0$. Next, we assume that $j_{k}\geq 1$. Again, Step 1 together with second equality in (23) yield $L_{k+1}:=2^{j_{k}-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}\geq 1$. To prove the second one, we assume by contradiction that $L_{k+1}>L+L_{0}$. This condition and (23) imply that $2^{j_{k}-1}L_{k}>L$. Thus, the first part of Proposition 5.1 implies that $j=j_{k}-1$ satisfies (22). Since the condition $L_{k+1}>L+L_{0}$ together (23) also implies $2^{j_{k}-1}L_{k}>L_{0}$, we have a contradiction because Proposition 5.1 also implies that $j_{k}$ is the smallest non-negative integer satisfying (27) and (28). Therefore, (34) holds for $k+1$, which concludes the proof that (34) holds for all $k\in{\mathbb{N}}$. The inequality (35) is an immediate consequence of (15), (21), the second inequality in (34) and (33). ∎
Next, we state iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1. Since the analysis closely follows that in [4], the proofs are omitted.
Theorem 5.3.
$\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is a stationary point to problem (14). In addition, for every $N\in{\mathbb{N}}$, there holds
$\min\left\{|{\omega}_{k}|:~k=0,1,\ldots,N\right\}\leq\max\left\{{{2(f(x^{0})-f^{*})}},\sqrt{2\alpha{(f(x^{0})-f^{*})}}\right\}\frac{1}{\sqrt{N+1}}.$
Since $\cal C$}} is compact and the functions $g$ and $h$ are convex, we can define the following constants:
$\Theta:=\min\left\{\frac{2}{{\bar{\Gamma}}\operatorname{diam}({\cal C})},\frac{1}{\alpha}\right\}\qquad{\bar{\Gamma}}:=\max_{x\in{{\cal C}}}\{\|\nabla g(x)\|+\|u\|:~u\in\partial h(x)\},\qquad\beta:=\frac{{\bar{\Gamma}}^{2}}{{\Theta L_{0}}}.$
Since the proof of the next theorem follows a standard argument in the convergence analysis of line–search methods, we omit it.
Theorem 5.4.
Let $\epsilon>0$. Then, Algorithm 1 generates a point $x^{k}$ such that $|{\omega}_{k}|\leq\epsilon$, performing at most,
$\left(\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}\right)\left(2+\log_{2}\left({\beta}\frac{1}{\epsilon^{2}}\right)\right)={\cal O}(|\log_{2}(\epsilon)|\epsilon^{-2})$
evaluations of functions $f$, and $\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}+1={\cal O}(\epsilon^{-2})$ evaluations of gradients of $f$ and subgradients of $h$.
5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
In this section, we establish iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 when the objective function is a DC function f=ghf=g-h that is piecewise weak–star–convex, where the first component $g$ has a Lipschitz continuous gradient on ${\cal C}\subset{\mathbb{R}}^{n}$ and the second component $h$ is convex but not necessarily differentiable. As will become clear, the piecewise-star-convexity property plays a central role in our analysis. By combining this property with the descent estimates from Proposition 5.1, we obtain the following global complexity result for the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1.
Theorem 5.5.
Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a piecewise-star-convex on $\cal C$}, and let ${\cal P}=\{\,{\cal C}_{i}:i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets such that $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. For each $i\in I$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
$f_{i}^{*}:=\min_{x\in{\cal C}_{i}}f(x)$
be the corresponding minimum value. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exists $i_{0}\in I$ such that:
$f(x^{k})-f_{i_{0}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k}$, for all $k=1,2,\ldots.$
$\min_{\ell\in\left\{\lfloor\frac{k}{2}\rfloor+2,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-2}$, for all $k=3,4,\ldots,$ where $\lfloor k/2\rfloor=\max_{n\in\mathbb{N}}\left\{n\leq k/2\right\}.$
Proof.
By (28) in Proposition 5.1, we have
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega(x^{k})|\,\lambda_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (36)
On the other hand, by (21),
$\lambda_{k}\;=\;\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big{\{}-|\omega(x^{k})|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big{\}}.$
Hence, for any choice $\beta_{k}\in(0,1]$, the right-hand side of (36) at $\lambda=\lambda_{k}$ is bounded by its value at $\lambda=\beta_{k}$. Taking $\beta_{k}:=\tfrac{2}{k+2}$, we get
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
Using $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$, together with (23) and Lemma 5.2 (which yields $\tfrac{2^{\,j_{k}}L_{k}}{2}\leq L+L_{0}$), we obtain
$f(x^{k+1})\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;(L+L_{0})\,\operatorname{diam}({\cal C})^{2}\,\beta_{k}^{2}.$ (37)
Since $\cal C$} is compact, the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1 is bounded. Let $\bar{x}$ be a limit point, and let $(x^{k_{\ell}})_{\ell\in\mathbb{N}}$ be a subsequence such that $\lim_{\ell\to+\infty}x^{k_{\ell}}={\bar{x}}$. Because ${\cal P}=\{{\cal C}_{i}:i\in I\}$ is a partition of $\cal C$}, there exists $i_{0}\in I$ with $\bar{x}\in{\cal C}_{i_{0}}$. In addition, Theorem 5.3 implies that $\lim_{k\to+\infty}{\omega}_{k}=0$ and $\bar{x}$} is a stationary point to problem (14). Hence, by (3.12), $\bar{x}$ is a minimizer of $f$ on ${\cal C}_{i_{0}}$. Let $f_{i_{0}}^{*}$ denote the minimum value of $f$ on ${\cal C}_{i_{0}}$. Thus, $\lim_{\ell\to+\infty}f(x^{k_{\ell}})=f({\bar{x}})=f_{i_{0}}^{*}$. Moreover, (29) in Proposition 5.1 implies that $\big{(}f(x^{k})\big{)}_{k\in\mathbb{N}}$ is monotone nonincreasing. Therefore,
$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{0}}^{*},\qquad\text{and}\qquad f_{i_{0}}^{*}\leq f(x^{k})\quad\forall\,k\in\mathbb{N}.$ (38)
Since $f$ is piecewise weak–star–convex on $\cal C$}, Proposition 3.11 and Theorem 2.2 ensure that, for any $x^{\star}\in X_{i_{0}}^{*}$,
$f_{i_{0}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$
where $u^{k}\in\partial h(x^{k})$. Using (20) and (26), the definition of $\omega_{k}$ gives $0\geq f_{i_{0}}^{*}-f(x^{k})\geq\omega_{k},$ hence
$0\leq f(x^{k})-f_{i_{0}}^{*}\leq|\omega_{k}|\qquad\forall\,k\in\mathbb{N}.$ (39)
Now apply Lemma 2.8 to the sequences
$a_{k}=f(x^{k})-f_{i_{0}}^{*}\leq b_{k}=|\omega_{k}|,\qquad\beta_{k}=\frac{2}{k+2},\qquad A=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$
The recursion (37) together with (39) satisfies the hypotheses of Lemma 2.8, yielding
$f(x^{k})-f_{i_{0}}^{*}\leq\frac{4(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k},\qquad\forall k\geq 1,$
and
$\min_{\ell\in\{\lfloor k/2\rfloor+2,\ldots,k\}}|\omega_{\ell}|\leq\frac{16(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k-2}.\qquad\forall k\geq 3.$
These are precisely items (i)–(ii) of Theorem 5.5. ∎
Theorem 5.5 shows that, under piecewise–star–convexity, the exact–gradient Frank–Wolfe method with our stepsize achieves the convex-case rate: $f(x_{k})-f^{\star}=\mathcal{O}(1/k)$ and the optimality certificate $|\omega_{k}|$ decays at the same order. In contrast, the next section focuses on the finite–difference variant. We do not pursue complexity bounds there; instead we establish asymptotic convergence under mild consistency assumptions on the difference radius and the adaptive stepsizes: the gradient surrogate converges to $\nabla g$ along the iterates, the scheme is well posed, and every accumulation point is first-order stationary (equivalently, the Frank–Wolfe gap and $|\omega_{k}|$ vanish). Thus the Frank–Wolfe method provides projection-free, asymptotic guarantees compatible with the piecewise–star–convex framework, even without claiming the convex-case $\mathcal{O}(1/k)$ rate.
6 Frank–Wolfe algorithm with finite difference
In this section we introduce a projection-free Frank–Wolfe variant that replaces $\nabla g$ by a finite-difference surrogate with an adaptive stepsize, while retaining the linear minimization oracle of Algorithm 1. The scheme is designed to the DC model f=ghf=g-h from earlier sections, $g$ smooth convex and $h$ convex (via subgradients), and fits the DC piecewise–star–convexity framework developed in previous sections. Our analysis targets asymptotic convergence of the finite-difference method: under mild consistency conditions on the probing radii and stepsizes, the surrogate tracks $\nabla g$, the algorithm is well posed, and every accumulation point is first-order stationary (the Frank–Wolfe gap and $|\omega_{k}|$ vanish). We do not claim rates for this variant; Theorem 5.5 remains the $\mathcal{O}(1/k)$ benchmark for the exact-gradient case. Since piecewise–star–convexity objectives are convex within cells, stabilization in a cell yields local convex geometry, which motivates our stepsize choice.
Let $\{e_{i}\}_{i=1}^{n}$ denote the canonical basis of $\mathbb{R}^{n}$. For a differentiable $g:\mathbb{R}^{n}\to\mathbb{R}$ and a stepsize $s>0$, define the (componentwise) forward finite-difference approximation of $\nabla g(x)$ by
$D(x,s):=\bigl{(}D_{1}(x,s),\ldots,D_{n}(x,s)\bigr{)},\qquad D_{i}(x,s):=\frac{g(x+se_{i})-g(x)}{s},\quad i=1,\ldots,n.$ (40)
The next lemma (see [38, (8.4), p. 195]) provides an upper bound on the error of approximating $\nabla g(x)$ by $D(x,s)$.
Lemma 6.1.
Assume that $g$ satisfies the assumption (A). Then, there holds
$\|\nabla g(x)-D(x,s)\|\leq\frac{{\sqrt{n}L}}{2}s,\qquad\forall x\in\mathbb{R}^{n},\quad\forall s>0.$
Proof.
Setting $\nabla g(x)=(\partial_{1}g(x),\ldots,\partial_{n}g(x))$ and using (40) we have
$\|\nabla g(x)-D(x,s)\|^{2}\leq\sum_{i=1}^{n}\left(\partial_{i}g(x)-D_{i}(x,s)\right)^{2}.$ (41)
Applying Proposition 2.4 with $v=e_{i}$ and $\lambda=s$, we conclude that $D_{i}(x,s)-\partial_{i}g(x)\leq(Ls)/2$. Thus, (41) implies that $\|\nabla g(x)-D(x,s)\|^{2}\leq n((Ls)/2)^{2}$, which implies the desired inequality. ∎
The statement of the Frank–Wolfe algorithm to solve problem (14) with finite difference approximating the gradient $\nabla g$ is as follows:
Algorithm 2.
Frank–WolfeC,f:=g-h algorithm with finite difference
Step 0.
Select $x^{0},x^{1}\in{{\cal C}}$, $x^{0}\neq x^{1}$ and $L_{1}>0$. Set $k=1$.
Step 1.
Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $i=0$.
Step 2.
Set
$s_{ij}=\frac{2L_{1}}{2^{i+j}L_{k}\sqrt{n}}\|x^{k}-x^{k-1}\|.$ (42)
Step 3.
Use an “LO oracle” to compute an optimal solution $p^{ij}$ and the optimal value ${\omega}_{ij}(x_{k})$ as follows
$p^{ij}\in{\rm argmin}_{p\in{{\cal C}}}(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{ij}(x_{k}):=(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{ij}-x^{k}).$ (43)
Step 4.
If ${\omega}_{ij}(x_{k})<0$, then set $i_{k}=i$ and go to Step 5. Otherwise, set $i\leftarrow i+1$, and go to Step 2.
Step 5.
Compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{i_{k}j}(x_{k})|}{{2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{i_{k}j}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda^{2}\right\}.$ (44)
Step 6.
If
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}-\frac{2^{j}L_{k}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$ (45)
then set $j_{k}=j$ and go to Step 7. Otherwise, set $j\leftarrow j+1$ and go to Step 2.
Step 7.
Set $\lambda_{k}:=\lambda_{j_{k}}$, $p^{k}:=p^{i_{k}j_{k}}$, define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
$x^{k+1}:=x^{k}+\lambda_{k}(p^{k}-x^{k}),\qquad\quad L_{k+1}:=2^{j_{k}-1}L_{k}.$ (46)
Set $k\leftarrow k+1$, and go to Step 1.
Let us describe the main features of Algorithm 2. First of all, note that due to $x^{0},x^{1}\in{\cal C}$, $p^{k}\in{\cal C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and $\cal C$} be convex set, it follows from inductive arguments and first equality in (46) that $(x^{k})_{k\in{\mathbb{R}}}\subset{\cal C}$. Note that (43) implies that
${\omega}_{ij}(x_{k})\leq D(x^{k},s_{ij})^{T}(p-x^{k}),\qquad p\in{\cal C}.$ (47)
Hence, considering that $x^{k}\in{\cal C}$, letting $p=x^{k}$ in the last inequality we conclude that
${\omega}_{ij}(x_{k})\leq 0.$ (48)
In order to simplify the notations, from now on we will use the following two notations:
$s_{k}:=s_{i_{k}j_{k}},\qquad\qquad{\omega}_{k}:={\omega}_{i_{k}j_{k}}(x_{k}).$ (49)
Proposition 6.2.
Assume that in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$. Then,
$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0,\qquad\forall p\in{{\cal C}},$ (50)
i.e., $x^{k}$ is a stationary point to problem (14).
Proof.
If in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$, then it follows from (47) that
$0\leq(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall i\in{\mathbb{N}},\quad\forall p\in{{\cal C}}.$ (51)
Since (42) implies that $\lim_{i\to+\infty}s_{ij}=0$, it follows from (40) that $\lim_{i\to+\infty}D(x^{k},s_{ij})=\nabla g(x^{k})$. Hence, taking limit in (51) we conclude that (50) holds. ∎
In view of Proposition 6.2, (48) and (49) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$.
6.1 Well-definedness
Herein, we establish that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined.
Lemma 6.3.
Let $s_{ij}$ be define in (42) and assume that ${2^{j}L_{k}}\geq 2L$. Then, there holds
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{L_{1}}}{2}\|x^{k}-x^{k-1}\|.$
Proof.
By applying Lemma 6.1 with $x=x^{k}$ and $s=s_{ij}$ we obtain that
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{\sqrt{n}L}}{2}s_{ij}$
Therefore, using (42) and considering that ${2^{j}L_{k}}\geq 2L$ the desired inequality follows. ∎
Lemma 6.4.
Let $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$. Then, (45) holds and $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying the following two conditions $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined and satisfies the following inequality:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$ (52)
Proof.
It follows from Proposition 2.4 that
$g(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (53)
On the other hand, considering that $u^{k}\in\partial h(x^{k})$ it follows from Proposition (2.5) that
$h(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{i_{k}j}-x^{k}\rangle$ (54)
Thus, combining (53) with (54) and taking into account the equality in (43) and (48), some algebraic manipulations shows that
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (55)
To proceed with the proof we need first to prove that for $\lambda_{j}$ defined in (44) we have
$|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}\leq-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}.$ (56)
For that we separately consider two cases: $\lambda_{j}=|{\omega}_{i_{k}j}(x_{k})|/{(2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2})}$ and $\lambda_{j}=1$. In the former, we have $2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2}=|{\omega}_{i_{k}j}(x_{k})|/{\lambda_{j}}$,which substituting into the left hand side of (56) yields an equality. If now $\lambda_{j}=1$, then it follows from (44) that $\lambda_{j}=1\leq|{\omega}_{i_{k}j}(x_{k})|/({2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2})$ or equivalently ${2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}\leq|{\omega}_{i_{k}j}(x_{k})|$ and $\lambda_{j}=1$, which again substituting into the left hand side of (56) yields the inequality. Hence, (56) holds. Therefore, (55) becomes
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (57)
Since ${2^{j}L_{k}}\geq 2(L+L_{1}/2)>L$, we can apply Lemma 6.3 to obtain that
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}.$
Hence, considering that $2\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}\leq\|x^{k}-x^{k-1}\|^{2}+\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}$, we have
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Thus, combining the last inequality with (57) and considering that $\lambda_{j}<1$, we conclude that
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{L+L_{1}/2-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Considering that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$, the last inequality implies that (45) holds. Since the first trial $j$ in (45) is defined in Step 1, i.e., $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $2^{j+1}L_{k}\geq 2^{j}L_{k}$ we conclude that $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, $(x^{k})_{k\in{\mathbb{N}}}$ is well defined, which proof the first three statements of the lemma. Therefore, (52) follows from the definition of $j_{k}$ in Step 6, (45) and (46), which concludes the proof. ∎
Since the proof of the next lemma is similar to Lemma 5.2 it will be omitted.
Lemma 6.5.
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{1}\leq L_{k}\leq 2(L+L_{1}),\quad\qquad k=1,2,\ldots.$ (58)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}}\right\},\quad\qquad k=1,2,\ldots.$ (59)
6.2 Convergence analysis
In this section we analyze the convergence properties of the sequence $(x^{k})_{k\in{\mathbb{N}}}$. For that, we assume that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is infinite.
Lemma 6.6.
The sequence $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}}$ is monotone and non-increasing. Moreover, there holds:
$\lim_{k\to+\infty}\|x^{k}-x^{k-1}\|=0.$ (60)
As a consequence, $\lim_{k\to+\infty}f(x^{k})={\bar{f}}$ for some ${\bar{f}}\geq f^{*}$.
Proof.
Since $L_{k}\geq L_{1}$, for all $k\in{\mathbb{N}}$, it follows from Lemma 6.4 that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$
which, taking into account that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, yields
$f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\leq f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots..$
which implies the first statement of the lemma. We proceed to prove the second statement. For that we first note that Lemma 6.5 implies that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$. Thus, using again that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, it follows from (52) that
$\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\leq f(x^{k})-f(x^{k+1}),\qquad k=1,2,\ldots.$ (61)
Let $\ell\in{\mathbb{N}}$ be such that $\ell>1$. Summing up the inequality (61) from $1$ to $\ell$ we obtain
$\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}\leq f(x^{1})-f^{*}.$ (62)
Considering that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$, some algebraic manipulations show that
$\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}+\frac{L_{1}}{4}\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}.$
Combining (62) with the last inequality we obtain that
$\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\frac{4}{L_{1}}\Big{(}f(x^{1})-f^{*}+\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}\Big{)}.$
Since the last inequality holds for all $\ell\in{\mathbb{N}}$, we have $\sum_{k=1}^{+\infty}\|x^{k+1}-x^{k}\|^{2}\leq+\infty$. Thus, (60) holds. To prove the last statement, we first note that due to $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in\mathbb{N}}$ be a non-increasing sequence and bounded from below by $f^{*}$, it must be convergent. Thus, we set ${\bar{f}}=\lim_{k\to+\infty}\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)$. Hence, considering that
$f(x^{k})=f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}-({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\geq f^{*},$
by using (60) we conclude that $\lim_{k\to+\infty}f(x^{k})={\bar{f}}\geq f^{*}$ and the proof is concluded. ∎
For simplifying the notations we define the following constants
${\bar{\alpha}}:=2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}>0.$ (63)
Lemma 6.7.
The following inequality holds:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},$ (64)
for all $k=1,2,\ldots$.
Proof.
Combining inequalities (52) with (58) and (59), and also using (63), we conclude that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\min\left\{1,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$
for all $k=1,2,\ldots$. Thus, (64) is an immediate consequence of the previous inequality. ∎
In the next theorem we will use the following constant
$\Gamma:=2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}.$ (65)
Theorem 6.8.
The following statements hold:
  • (i)
    $\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is a stationary point to problem (14);
  • (ii)
    for every $N\in{\mathbb{N}}$, there holds
    $\min\left\{|{\omega}_{k}|:~k=1,\ldots,N\right\}\leq\max\left\{{\Gamma},\sqrt{{\bar{\alpha}}\Gamma}\right\}\frac{1}{\sqrt{N}}.$
Proof.
To prove the item (i), we first note that (64) is equivalent to
$\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2}\Big{)}-\Big{(}f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\Big{)},$ (66)
for all $k=1,2,\ldots$. It follows from Lemma 6.6 that $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}^{*}}$ monotone and non-increasing. Thus, considering that it is also bounded from below by $f^{*}$, we conclude that it converges. Hence, taking limit in (66) we have
$\lim_{k\to+\infty}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}=0,$
which implies the first statement of item (i). For proving the second statement of item (i), take $\bar{x}$} a limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ and subsequence $(x^{k_{i}})_{i\in{\mathbb{N}}}$ such that $\lim_{i\to+\infty}x^{k_{i}}={\bar{x}}$. Since $\cal C$}} is a compact set the sequence $(u^{k})_{k\in{\mathbb{N}}^{*}}$ is bounded. Thus, without loss of generality we assume that $\lim_{i\to+\infty}u^{k_{i}}={\bar{u}}$. On the other hand, Lemma 6.6 together with (42) imply that
$\lim_{k\to+\infty}s_{k}=0.$ (67)
Let $(s_{k_{i}})_{i\in{\mathbb{N}}^{*}}$ subsequence of $(s_{k})_{k\in{\mathbb{N}}^{*}}$. Thus, we conclude from (40) that
$\lim_{i\to+\infty}D(x^{k_{i}},s_{k_{i}})=\nabla g({\bar{x}}).$ (68)
Finally, using (43) we conclude that
$(D(x^{k_{i}},s_{k_{i}})-u^{k_{i}})^{T}(p-x^{k_{i}})\geq{\omega}_{k_{i}},\qquad\forall p\in{{\cal C}}.$
Therefore, taking limite in the last inequality and using (67) together (68) we obtain that
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(p-{\bar{x}})\geq 0,\qquad{\bar{u}}\in\partial g({\bar{x}}),\qquad\forall~x\in{{\cal C}},$
which implies that $\bar{x}$} is a stationary point to problem (14). To prove the item (ii), take $N\in{\mathbb{N}}^{*}$. Thus, (66) implies that
$\sum_{k=1}^{N}\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{1})+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\Big{)}-\Big{(}f(x^{N+1})+\frac{L_{1}}{2}\|x^{N}-x^{N-1}\|^{2}\Big{)}.$
Hence, we conclude that
$\min\left\{\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}:~k=1,\ldots N\right\}\leq\frac{1}{N}\left(f(x^{1})-f^{*}+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\right).$
It follows from the last inequality that there exists ${\bar{k}}\in\{1,\ldots N\}$ such that
$|{\omega}_{\bar{k}}|\leq\frac{1}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right),\qquad|{\omega}_{\bar{k}}|\leq\sqrt{\frac{{\bar{\alpha}}}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right)},$
which, by using (65), implies the item (ii). ∎
7 Conclusions
We analyzed Algorithm 1 and its finite–difference variant Algorithm 2 within the DC piecewise-star-convex framework. For piecewise-star-convex objectives, the exact–gradient method attains the convex–case rate $\mathcal{O}(1/k)$ for both the function suboptimality $f(x_{k})-f^{\star}$ and the optimality certificate (duality/Frank–Wolfe gap). For the finite–difference method, we established asymptotic convergence and an $\mathcal{O}(1/\sqrt{k})$ decay of the gap. A full iteration–complexity theory for the finite–difference variant under piecewise-star-convex remains open. Promising directions include: (i) matching $\mathcal{O}(1/k)$ rates via tighter control of finite–difference bias and adaptive radii, (ii) guarantees that account for possible transitions across piecewise-star-convex cells without stabilization assumptions, and (iii) extensions to stochastic oracles, away–steps, and line–search variants.
Funding
The first and third authors were supported by the Australian Research Council (ARC), Solving hard Chebyshev approximation problems through nonsmooth analysis (Discovery Project DP180100602). The second author was supported in part by CNPq - Brazil Grants 304666/2021-1 and Fapesc - TR nº 2024TR002238. This work was partially conducted while the second author visited the first and third authors at Deakin University in November 2022, with thanks to the host institution for funding and providing a stimulating scientific environment.

### v3

@@
-Frank–Wolfe algorithm for DC optimization problem
+Frank–Wolfe algorithms for piecewise star-convex functions with a nonsmooth difference-of-convex structure
 Abstract
-We propose two versions of the Frank–Wolfe (conditional gradient) algorithm with adaptive step sizes for DC optimization problems, where the objective is the difference of a differentiable convex function with Lipschitz continuous gradient and a convex (possibly non-differentiable) function.
-The second version builds on the first, using finite differences to approximate the gradient of the first component.
-Unlike previous approaches relying on curvature or Lipschitz constants, our stepsizes require no such parameters.
-We also introduce the class of piecewise-star-convex functions, which may be nonconvex and non–differentiable, and show that the first version achieves an $1/k$ convergence rate for minimizing such functions, while ensuring stationarity of all limit points.
-In the second version, the finite–difference scheme yields relative rather than absolute gradient errors, simplifying the analysis; we prove that all limit points are stationary and that the duality gap decreases at rate $1/\sqrt{k}$.
-Keywords: Frank-Wolfe method; DC optimization problem; finite difference; weak-star-convex function.
+We study Frank–Wolfe algorithms for minimizing a nonconvex function expressed as the difference of a differentiable convex function with a Lipschitz continuous gradient and a possibly nonsmooth convex function over a compact convex set.
+We propose an adaptive Frank–Wolfe scheme that exploits this difference-of-convex structure and prove a sufficient descent property together with the convergence of the Frank–Wolfe gap to zero.
+We also introduce the notion of piecewise star-convexity, and show that for this class of functions Clarke-stationary points are minimizers on the cells of the underlying partition and, for finite partitions, we establish $\mathcal{O}(1/k)$ complexity bounds in function value and in the Frank–Wolfe gap.
+In particular, these bounds match the standard rates obtained in the convex case, showing that Frank–Wolfe methods can retain optimal complexity guarantees beyond the classical convex setting, even in the presence of nonsmooth terms.
+Keywords: Frank-Wolfe algorithm; DC optimization problem; finite difference; piecewise-star-convex function.
 AMS subject classification: 90C25, 90C60, 90C30, 65K05.
@@
 The DC optimization method involves minimizing a function that can be represented as the difference between two convex functions.
-We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable.
+We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ are continuously differentiable convex functions, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable.
 To the best of our knowledge, DC optimization can be traced back to pioneering works such as [37, 38], which introduced the first algorithms for this problem.
@@
 Since then, this method has attracted the attention of several researchers who work with continuous optimization, thus becoming also known as the conditional gradient method.
-One of the factors that explain the interest in this method is its simplicity and ease of implementation.
+One of the factors that explains the interest in this method is its simplicity and ease of implementation.
 In fact, each method iteration only requires access to a linear minimization oracle over a compact convex set.
@@
 For all of these reasons, several variants of this method have arisen throughout the years and new properties of it have been discovered, resulting in a large literature on it, papers dealing with this method include [4, 9, 10, 16, 17, 19, 24, 27, 31].
-In the present paper, we formulate two versions of Frank–Wolfe algorithm or conditional gradient method to solve DC optimization problem with an adaptive step size.
-The second version is based on the first and employs finite differences to approximate the gradient $\nabla g$.
+In the present paper, we formulate a version of Frank–Wolfe algorithm or the conditional gradient method to solve DC optimization problem with an adaptive step size.
 It is worth mentioning that the Frank-Wolfe algorithm has previously been formulated in the context of DC programming, see [23].
@@
 In contrast to the previous study, which assume that the curvature/Lipschitz-type constant of f=ghf=g-h is bounded from above, the analysis of the Frank-Wolfe method done here just assumes that the gradient of the first component $g$ of $f$ is Lipschitz continuous.
-While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous in both formulations, we will not compute the step size using the Lipschitz constant of $\nabla g$, as in earlier formulations that employed the curvature/Lipschitz-type constant of the objective function $f$.
+While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous, we will not compute the step size using the Lipschitz constant of $\nabla g$.
 The step size will be computed adaptively, based on an idea introduced in [3] (see also [5, 35]) that approximates the Lipschitz constant.
 It has been shown in previous works that Frank-Wolfe algorithm produces a stationary point with a convergence rate of only ${\cal O}(1/\sqrt{k})$ for non-convex objective functions, see [26] (see also [23]).
-We introduce the concept of weak-star-convex function notion, which generalizes the star-convex function concept proposed in [37] as well as a few other related concepts, such as [22, 45].
-The weak-star-convex objective functions are also taken into account in the convergence analysis of the first method presented here.
-The convergence rate is $\mathcal{O}(1/k)$ for both the function values and the duality gap, even though piecewise-star-convexs are generally non-differentiable and non-convex.
-Consequently, the class of functions for which the Frank–Wolfe method achieves an $\mathcal{O}(1/k)$ convergence rate includes piecewise-star-convexs that are differences of convex functions, with the first component having a Lipschitz continuous gradient.
-This class complements the set of star–convex functions, which was recently shown in [35] to enjoy the same convergence rate.
-As mentioned previously, the second version of the Frank-Wolfe algorithm uses finite difference to approximate the gradient of the function $g$.
-Similar to the strategies adopted in [20], the finite difference utilized here to approximate the gradient is computed with the step-size updated adaptively utilizing two previous iterations.
-Finite difference is an old concept that has been used to approximate derivatives in optimization settings (for example, see [38, Section 8.1]) and has already appeared in the study of the Frank-Wolfe algorithm (for example, see [14, 16, 25, 40]).
-It is worth noting that all previous applications of finite difference in the Frank-Wolfe algorithm provided approximate gradients with absolute error, whereas the one used here has the advantage of providing us with a relative error and thus being quite simple to analyze.
-We show that all limit points of the generated sequence for the second version of the Frank-Wolfe algorithm are stationary points for the problem under consideration, and we establish that the rate of convergence is ${\cal O}(1/\sqrt{k})$ for the duality gap.
+Here, we introduce the concept of piecewise star-convex function notion, which generalizes the star-convex function concept proposed in [34] as well as a few other related concepts, such as [20, 40].
+The convergence analysis developed in this paper covers piecewise star–convex with the nonsmooth difference-of-convex structure aforementined, we show that the proposed Frank–Wolfe scheme equipped with our backtracking stepsize attains the same order of iteration complexity as in the convex case, both the function-value and the Frank-Wolfe gap decay at the standard sublinear rate $\mathcal{O}(1/k)$.
+This holds even though piecewise star–convex functions are, in general, nonconvex and may be nonsmooth across cell boundaries.
+Consequently, our theory extends convex-type complexity guarantees for Frank–Wolfe methods to a broader nonconvex setting, including piecewise star–convex DC objectives with a smooth convex term $g$ and a convex, possibly nonsmooth term $h$.
+These guarantees complement recent results for star–convex functions in [32], and indicate that projection-free Frank–Wolfe schemes can maintain the usual first-order complexity behavior beyond classical convexity, even in the presence of nonsmooth terms.
 The remainder of the paper is organized as follows.
 Section 2 introduces the notation and collects auxiliary results.
-In Section 3 we define the concept of piecewise–star–convex, present its main properties, and provide illustrative examples.
-Section 4 states the optimization problem, the required assumptions, and related notation.
-Section 5 is devoted to the first version of the Frank–Wolfe algorithm, including its formulation, well–posedness, and analysis of the generated sequence.
-Section 6 presents the second version of the algorithm together with its formulation, well–posedness, and convergence analysis.
+Section 3 presents the optimization problem, the standing assumptions, and related notation.
+In Section 4 we define piecewise star–convexity, establish its main properties, and provide illustrative examples.
+Section 5 is devoted to the proposed Frank–Wolfe algorithm, including its formulation, well-posedness, and the basic properties of the generated sequence.
+Section 5.1 analyzes the convergence properties of the method, and Section 5.2 derives the corresponding iteration–complexity bounds.
 Finally, Section 6 summarizes our conclusions.
 2 Preliminaries
-In this section, we recall some notations, definitions and basics results used throughout the paper.
+In this section, we recall some notations, definitions and basic results used throughout the paper.
 A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be convex if $\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)$, for all $x,y\in{\mathbb{R}^{n}}$ and $\lambda\in[0,1]$.
-And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex function see [23].
+And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex functions see [21].
 We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz if, for all $x\in\mathbb{R}^{n}$, there exist a constant ${\cal C}_{x}>0$ and a neighborhood $U_{x}$ of $x$ such that $|f(x)-f(y)|\leq{\cal C}_{x}\|x-y\|$, for all $y\in U_{x}.$ It is well known that, if $f:\mathbb{R}^{n}\to\mathbb{R}$ is convex, then $f$ is locally Lipschitz.
-And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [10, p.
+And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [11, p.
 32].
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function.
-The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial_{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
+The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial^{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
 $f^{\circ}(x;d)=\limsup_{\tiny{\begin{array}[]{c}u\rightarrow x\\ t\downarrow 0\end{array}}}\frac{f(u+td)-f(u)}{t}.$
-For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [10, p.
+For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [11, p.
 27].
-If $f$ is convex, then $\partial_{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [10, p.
+If $f$ is convex, then $\partial^{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [11, p.
 36].
-We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial_{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [10, p.
+We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial^{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [11, p.
 33].
-Theorem 2.1 ([10, p. 27]).
+Theorem 2.1 ([11, p. 27]).
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function.
-Then, $\partial_{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial_{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$.
-Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial_{c}f(x)\}$.
-Theorem 2.2 ([10, p. 38-39]).
+Then, $\partial^{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial^{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$.
+Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial^{c}f(x)\}$.
+Theorem 2.2 ([11, p. 38-39]).
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be given by f=ghf=g-h, where $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable.
-Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial_{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
-The next result is a combination of Theorem 2.2 with [10, Corollary on p. 52].
+Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial^{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
+The next result is a combination of Theorem 2.2 with [11, Corollary on p. 52].
 Theorem 2.3.
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function and ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a closed and convex set.
-If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and all $x\in{{\cal C}}$.
-As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $u\in\partial_{c}h({x^{*}})$ and all $x\in{{\cal C}}$.
-Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$.
+If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then there exists $v\in\partial^{c}f(x^{*})$ such that $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $x\in{{\cal C}}$.
+As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then there exists $u\in\partial^{c}h({x^{*}})$ such that $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, all $x\in{{\cal C}}$.
+Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for some $v\in\partial^{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$.
 A continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ has gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$ if there exists a Lipschitz constant $L>0$ such that $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
-Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [6, Proposition A.24], see also[12, Lemma 2.4.2].
+Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [7, Proposition A.24], see also[13, Lemma 2.4.2].
 Proposition 2.4.
@@
 The function $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex if and only if $h(y)\geq h(x)+\langle u,y-x\rangle$, for all $x,y\in\mathbb{R}^{n}$ and all $u\in\partial h(x)$.
-Proposition 2.6 ([23, Proposition 6.2.1]).
+Proposition 2.6 ([21, Proposition 6.2.1]).
 Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex.
-Let $(x^{k})_{k\in{\mathbb{N}}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$.
+Let $\{x^{k}\}_{k\in\mathbb{N}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$.
 If $\lim_{k\to+\infty}x^{k}={\bar{x}}$ and $\lim_{k\to+\infty}u^{k}={\bar{u}}$, then ${\bar{u}}\in\partial h({\bar{x}})$.
-Proposition 2.7 ([23, Proposition 6.2.2]).
+Proposition 2.7 ([21, Proposition 6.2.2]).
 Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex.
@@
 the image $\partial h(B)$ of a bounded set $B\subset\mathbb{R}^{n}$ is a bounded set in $\mathbb{R}^{n}$.
-In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [2, Lemma 13.13, Ch. 13, p. 387].
+In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [6, Lemma 13.13, Ch. 13, p. 387].
 Lemma 2.8.
@@
 In particular, the $\ell_{1}$–norm is $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, and the $\ell_{\infty}$–norm is $\|x\|_{\infty}:=\max_{1\leq i\leq n}|x_{i}|$.
-3 Piecewise-star-convex functions
-In this section, we introduce the concept of piecewise-star-convexity, a relaxation of the classical star–convexity notion from [37].
-Many DC functions that are nonconvex nevertheless satisfy this weaker property.
-As will be shown in a later section, by identifying and exploiting piecewise–star–convexity we can broaden the class of problems for which the classical Frank–Wolfe algorithm attains an ${\cal O}(1/k)$ convergence rate.
-We begin by recalling the definition of star–convexity from [37] and then present its relaxed counterpart, the piecewise-star-convex notion, which will be central to our analysis.
+3 The DC optimization problem
+We are interested in solving the following constrained DC optimization problem
+$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (1)
+where ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact and convex set, $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function and $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex function possibly non-differentiable, with domain on $\mathbb{R}^{n}$.
+Throughout the paper we assume that the gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, i.e., there exists a Lipschitz constant $L>0$ such that
+  • (A)
+    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
+Since we are assuming that ${\cal C}\subset{\mathbb{R}}^{n}$ is a compact set, its diameter is a finite number defined by
+$\operatorname{diam}({\cal C}):=\max\left\{\|x-y\|:~x,y\in{\cal C}\right\}.$ (2)
+Since $\cal C$} is compact and $f$ is continuous on $\cal C$}, the problem (1) admits a finite global optimal value
+$f_{\cal C}^{*}:=\min_{x\in{\cal C}}f(x)\;>\;-\infty,$ (3)
+and the set of minimizers ${\cal C}^{*}:=\arg\min_{x\in{\cal C}}f(x)$ is nonempty.
+By Theorem 2.3, a point $\bar{x}\in{\cal C}$ is said to be Clarke stationary for problem (1) if there exists $\bar{u}\in\partial^{c}h(\bar{x})$ such that
+$\bigl(\nabla g(\bar{x})-\bar{u}\bigr)^{\top}(x-\bar{x})\;\geq\;0,\qquad\forall\,x\in{\cal C}.$ (4)
+In general, (4) is a necessary condition for optimality but may fail to be sufficient. In particular, every minimizer $x^{*}\in{\cal C}^{*}$ satisfies (4).
+Let us introduce the linear minimization oracle and the associated gap function that will be used to define the Frank–Wolfe step for the DC model (1).
+Since f=ghf=g-h may be nonsmooth, the linearization is built from $\nabla g(x)$ and a subgradient of $h$.
+Given $x\in{\cal C}$ and $u\in\partial h(x)$, define
+$p(x,u)\in\arg\min_{v\in{\cal C}}\bigl(\nabla g(x)-u\bigr)^{\top}(v-x),\qquad\omega(x,u):=\bigl(\nabla g(x)-u\bigr)^{\top}\bigl(p(x,u)-x\bigr).$ (5)
+When no ambiguity arises, we write $p(x)$ and $\omega(x)$ for $p(x,u)$ and $\omega(x,u)$, where $u$ is the algorithmic choice at $x$.
+Proposition 3.1.
+Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex.
+For every $x\in{\cal C}$ and every $u\in\partial^{c}h(x)$, let $p(x,u)$ and $\omega(x,u)$ be defined by (5).
+Then $\omega(x,u)\leq 0.$ Moreover, if $\omega(x,u)=0$, then $x$ is a stationary point for (1).
+Conversely, if $x\in{\cal C}$ is stationary, then there exists $u\in\partial^{c}h(x)$ such that $\omega(x,u)=0$.
+Proof.
+Fix $x\in{\cal C}$ and $u\in\partial^{c}h(x)$. Since the point $v=x$ is feasible in (5), we have
+$\omega(x,u)=\min_{v\in{\cal C}}\bigl(\nabla g(x)-u\bigr)^{\top}(v-x)\leq\bigl(\nabla g(x)-u\bigr)^{\top}(x-x)=0,$
+which proves $\omega(x,u)\leq 0$.
+Furthermore, by (4), there exists $\bar{u}\in\partial^{c}h(\bar{x})$ such that $w(x,u)\geq 0$ if and only if $x$ is Clarke stationary, which completes the proof. ∎
+We conclude this section with a standard example of a nonsmooth DC function, see a particular instance in [1, Example 2.4.].
+It satisfies our standing assumptions.
+In particular, it admits a decomposition f=ghf=g-h in which $g$ is convex, differentiable, and has a Lipschitz continuous gradient, while $h$ is convex and possibly nonsmooth.
+Example 3.2.
+Let $A\in\mathbb{R}^{n\times n}$ have full column rank, so that $Q:=A^{\top}A$ is positive definite. Let $b\in\mathbb{R}^{n}$ and $\beta>0$. Define $f:{\mathbb{R}}^{n}\to\mathbb{R}$ by
+$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{2}-\beta\|x\|_{1}.$
+Then $f$ is a nonsmooth DC function with the decomposition f=ghf=g-h, where $g(x):=\tfrac{1}{2}\,\|Ax-b\|_{2}^{2}$ and $h(x):=\beta\|x\|_{1}$.
+Moreover, $g$ is convex and differentiable with a Lipschitz continuous gradient, while $h$ is convex and nonsmooth.
+Beyond its DC structure, this example illustrates a geometric feature that will be central in the sequel.
+Although $f$ is nonconvex, it becomes well behaved when restricted to suitable convex regions.
+Indeed, the sign pattern of $x$ induces a natural partition of $\mathbb{R}^{n}$ into orthants, and on each orthant the function reduces to a strictly convex quadratic.
+Consequently, any Clarke-stationary point is a minimizer of $f$ on its corresponding cell, even if it is not a global minimizer on $\mathbb{R}^{n}$.
+This motivates the class of piecewise star-convex models introduced in the next section, where star-type inequalities hold on each cell of a convex partition and can be used to obtain sharper convergence and complexity guarantees for Frank–Wolfe methods.
+4 Piecewise star-convexity with a nonsmooth DC structure
+In this section, we introduce the notion of piecewise star-convexity, a relaxation of the classical concept of star convexity from [34].
+Although a difference-of-convex function is generally nonconvex, it often exhibits this weaker geometric property when restricted to suitable regions of the domain.
+As we show later, identifying and exploiting piecewise star-convexity yields sublinear convergence guarantees for Frank–Wolfe methods.
+Moreover, when the domain admits a finite partition, we obtain an ${\cal O}(1/k)$ rate on each region that matches the standard rate of the convex case, despite moving beyond the classical convex setting and allowing nonsmooth terms.
+We begin by recalling the definition of star convexity from [34] and then introduce its piecewise counterpart, which will play a central role in our analysis.
 Definition 4.1.
@@
 Indeed, letting the following convex functions $g(s,t)=(s^{2}+t^{2})^{2}+s^{2}+t^{2}$ and $h(s,t)=s^{2}t^{2}+s^{4}+t^{4}$, we have f=ghf=g-h.
-The requirement in (1) is formulated with respect to a fixed minimizer $x^{*}$ and must hold for all $x\in{\cal C}$.
-A milder, yet still powerful, relaxation is to allow the reference minimizer to depend on $x$.
-In many DC models, the reference minimizer that makes (1) valid for a given $x$ may vary with $x$, which is precisely what piecewise-star-convexity captures.
-Even this global condition can be too restrictive when $f$ exhibits multiple “basins” of favorable geometry.
-A natural remedy is to partition the domain into regions within which a local star–type inequality holds with respect to a (local) minimizer:
+The condition in (6) is stated with respect to a fixed minimizer $x^{\ast}$ and must hold for all $x\in{\cal C}$.
+A milder relaxation is to allow the reference minimizer to depend on $x$, a behavior that often arises in DC models and is captured by piecewise star-convexity.
+When $f$ has several “basins” of favorable geometry, even this can be too restrictive globally, so we partition the domain into regions where a star-type inequality holds relative to a local minimizer:
 Definition 4.3.
-Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets with pairwise disjoint interiors and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$.
+Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$.
 We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$} if, for every $i\in I$,
@@
 Remark 1.
-When the partition is trivial (${\cal P}=\{{\cal C}\}$) and $X_{1}^{*}=X^{*}$, Definition 3.3 reduces to Definition 3.1.
-In practice, piecewise-star-convexity often arises when $f$ is convex on each region ${\cal C}_{i}$ but not across different regions.
-In this setting, the minimizers $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ are global minimizers within their corresponding cells ${\cal C}_{i}$, although they may be only local (and possibly non–global) minimizers of $f$ on the whole domain $\cal C$}.
-This structure ensures that star–type inequalities hold on each cell (with respect to its minimizers $X_{i}^{*}$), even if such inequalities fail when crossing cell boundaries.
-We emphasize that in Definition 3.3, the collection of subsets in item (ii) need not be finite: the piecewise-star-convexity property may still hold when $\cal C$} is covered by infinitely many convex regions.
-The following example illustrates that finiteness cannot, in general, be assumed.
+When the partition is trivial, that is, ${\cal P}=\{{\cal C}\}$ and $X_{1}^{*}=X^{*}$, Definition 4.3 reduces to Definition 4.1.
+In many applications, piecewise star-convexity arises because $f$ has a favorable geometry on each region ${\cal C}_{i}$ (for instance, it may be convex or star-convex when restricted to ${\cal C}_{i}$) but this geometry does not persist across different regions.
+Accordingly, the sets $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ consist of global minimizers of the cellwise problems, yet such points need not be global minimizers of $f$ on $\cal C$} and may fail to be even local minimizers of $f$ on $\cal C$}.
+The key point is that the star-type inequalities in Definition 4.3 are enforced within each cell relative to $X_{i}^{*}$, while no compatibility is required when one crosses cell boundaries.
+We emphasize that, in Definition 4.3, the collection of subsets in item (ii) need not be finite.
+Piecewise star-convexity may still hold when $\cal C$} is covered by infinitely many convex regions.
+The following example exhibits an infinite partition and an objective function with a nonsmooth difference-of-convex structure.
 Example 4.4.
@@
 $V_{1}=\Bigl[\tfrac{3}{4},\,1\Bigr],\quad V_{k}=\Bigl[\tfrac{1/k+1/(k+1)}{2},\ \tfrac{1/k+1/(k-1)}{2}\Bigr],\ k\geq 2,\quad V_{0}=\{0\}.$
-By construction, $x\in V_{x_{x}^{*}}$ and the whole segment $x_{x}^{*},x]$ lies in $V_{x_{x}^{*}}$. Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y\in[x_{x}^{*},x]$. Hence, for any $\lambda\in[0,1]$,
+By construction, $x\in V_{x_{x}^{*}}$ and the whole segment between $x_{x}^{*}$ and $x$ lies in $V_{x_{x}^{*}}$.
+Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y$ belonging the segment between $x_{x}^{*}$ and $x$.
+Hence, for any $\lambda\in[0,1]$,
 $f\big(\lambda x_{x}^{*}+(1-\lambda)x\big)=\big|\lambda x_{x}^{*}+(1-\lambda)x-x_{x}^{*}\big|^{2}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x).$
@@
 The black segments indicate the active arc of the lower envelope, the dashed lines mark the boundaries of $V(x_{x}^{*})$, and the solid points on the $x$-axis represent the centers $c\in S$.
-We next present a prototypical example of a function that is piecewise-star-convex.
-This model arises naturally in sparse estimation, robust quadratics, and regularized control.
-The $\ell_{1}$ penalty acts as a sign–dependent linear perturbation of a positive–definite quadratic.
-By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the function reduces on each region to a strictly convex quadratic with a unique minimizer.
-These cell minimizers provide the reference points in Definition 3.3, thereby establishing piecewise-star-convexity even though, globally, such minimizers may be only local.
+We next present a prototypical example of a piecewise star-convex function with a nonsmooth DC structure.
+This model arises naturally in sparse estimation, robust quadratic fitting, and regularized control.
+The $\ell_{1}$ term is convex but nonsmooth, and it induces a sign-dependent linear perturbation of a positive-definite quadratic.
+By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the objective reduces on each region to a strictly convex quadratic with a unique minimizer.
+These cell minimizers provide the reference points in Definition 4.3, which establishes piecewise star-convexity even though, globally, the problem remains nonconvex.
 For use in the examples below and to simplify notation, define
@@
 For each $s\in\{-1,1\}^{n}$, define the closed orthant
-$\Omega_{s}=\bigl{\{}\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0,\,\text{for }i=1,\ldots,n\bigr{\}},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
-On each $\Omega_{s}$, the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$, for $x\in\Omega_{s}$.
-Hence, the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
-$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$
+$\Omega_{s}=\bigl\{\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0\text{ for }i=1,\ldots,n\bigr\},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
+On each $\Omega_{s}$ the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$ for $x\in\Omega_{s}$.
+Hence the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
+$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$ (10)
 with constant Hessian $Q$.
-Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\mathbb{R}^{n}$ with modulus $\lambda_{\min}$, the minimizer $x_{s}^{*}$ is unique.
+Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\Omega_{s}$ with modulus $\lambda_{\min}(Q)$; hence the minimizer $x_{s}^{*}$ is unique.
 Considering that $f$ coincides with the strictly convex quadratic $f_{s}$ on each orthant $\Omega_{s}$ and $x_{s}^{*}$ is the (unique) minimizer of $f$ on $\Omega_{s}$, the convexity inequality
 $f\bigl(\lambda x_{s}^{*}+(1-\lambda)x\bigr)\leq\lambda f(x_{s}^{*})+(1-\lambda)f(x),\qquad\forall\,x\in\Omega_{s},\ \forall\,\lambda\in[0,1],$
-holds. Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$.
-Next we preste a constrained DC optimization problem motivated by contrastive learning formulations.
-Example 3.6.
-Given two disjoint sets of reference points, a set of positive examples $\mathcal{P}=\{\bar{p}_{1},\dots,\bar{p}_{m}\}\subset\mathbb{R}^{n}$ and a set of negative examples $\mathcal{N}=\{\bar{q}_{1},\dots,\bar{q}_{r}\}\subset\mathbb{R}^{n}$, the objective is to identify a point $x$ that lies close to the positive references while remaining far from the negative ones.
-Here all distances are measured with respect to the standard Euclidean metric.
-To remain consistent with our general framework, we impose that the search is carried out over a closed and convex feasible set $\mathcal{C}\subset\mathbb{R}^{n}$.
-Formally, we study the problem
-$\min_{x\in\mathcal{C}}f(x),$ (5)
-where the Euclidean contrastive objective is defined by
-$f(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}-\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},\qquad x\in\mathbb{R}^{n},$ (6)
-with $\lambda_{i}^{+}>0$ and $\lambda_{j}^{-}>0$ fixed weights.
-The squared $\ell_{2}$ terms pull $x$ toward the positive examples, while the negative $\ell_{1}$ term acts as a contrastive regularizer, promoting separation from the negative examples in a coordinate-dependent way.
-The objective (6) admits a natural DC decomposition of the form
-$g(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2},\qquad h(x):=\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},$ (7)
-where $g$ is smooth and strongly convex, while $h$ is convex but nondifferentiable.
-Hence similarly to Example 3.5 we can shoe that f=ghf=g-h belongs to the class of piecewise-star-convex DC models.
-In fact, let $m\geq 1$ and $\lambda_{i}^{+}>0$ for $i=1,\ldots,m$, and define $\Lambda^{+}:=\sum_{i=1}^{m}\lambda_{i}^{+}>0$.
-Consider $f:\mathbb{R}^{n}\to\mathbb{R}$ given in (6).
-Then $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the partition ${\cal P}=\{{\cal C}_{S}\}_{S}$ induced by the hyperplane arrangement $\{x_{k}=\bar{q}_{j,k}:~j=1,\ldots,r,\ k=1,\ldots,n\}$, where for $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$,
-${\cal C}_{S}:=\bigl{\{}\,x\in\mathbb{R}^{n}:\ s_{j,k}\,(x_{k}-\bar{q}_{j,k})\geq 0,\ \forall j=1,\ldots,r,\ \forall k=1,\ldots,n\,\bigr{\}}.$
-Fix $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$ and consider $x\in{\cal C}_{S}$. For every $j$ and $k$ we have $|x_{k}-\bar{q}_{j,k}|=s_{j,k}(x_{k}-\bar{q}_{j,k})$ on ${\cal C}_{S}$, hence
-$\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1}\;=\;\sum_{j=1}^{r}\lambda_{j}^{-}\sum_{k=1}^{n}s_{j,k}(x_{k}-\bar{q}_{j,k})\;=\;v_{S}^{\top}x-c_{S},$
-where $v_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}\in\mathbb{R}^{n}$ and $c_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}^{\top}\bar{q}_{j}$ is a constant on ${\cal C}_{S}$. Moreover,
-$\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}\;=\;\Lambda^{+}\|x\|_{2}^{2}-2\,b^{\top}x\;+\;d,$
-with $b:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\bar{p}_{i}$ and $d:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|\bar{p}_{i}\|_{2}^{2}$.
-Therefore, on ${\cal C}_{S}$ the function $f$ reduces to the strongly convex quadratic
-$f_{S}(x)\;=\;\Lambda^{+}\|x\|_{2}^{2}-(2b+v_{S})^{\top}x\;+\;(d-c_{S}),$
-whose Hessian is $2\Lambda^{+}I$ and is positive definite.
-Since ${\cal C}_{S}$ is closed and convex and $f_{S}$ is strongly convex, $f_{S}$ admits a unique minimizer $x_{S}^{\ast}\in{\cal C}_{S}$.
-By convexity of $f_{S}$ on the convex set ${\cal C}_{S}$, for all $x\in{\cal C}_{S}$ and all $\lambda\in[0,1]$,
-$f\bigl{(}\lambda x_{S}^{\ast}+(1-\lambda)x\bigr{)}\;\leq\;\lambda f(x_{S}^{\ast})+(1-\lambda)f(x).$
-This is precisely the star–type inequality required in Definition 3.3 with anchor $x_{S}^{\ast}$ on the cell ${\cal C}_{S}$.
-Since the interiors of the cells $\{{\cal C}_{S}\}$ are pairwise disjoint and $\bigcup_{S}{\cal C}_{S}=\mathbb{R}^{n}$, the claim follows.
-The structure f=ghf=g-h reflects the central principle of contrastive learning: the quadratic component promotes similarity with positive references, while the $\ell_{1}$ component enforces dissimilarity from negative ones.
-For comprehensive discussions of contrastive learning in the Euclidean setting, see [24].
-In the next proposition, we establish that, under a suitable dominance condition on the positive weights, the function $f$ in (6) is coercive and therefore attains at least one global minimizer of (5).
-The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization (see [Libertietall2014] and [1] for a comprehensive survey on this subject).
+holds.
+Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$.
+A particularly interesting case is when $Q=\alpha I_{n}$ with $\alpha>0$.
+In this case, it follows from (10) that on each orthant $\Omega_{s}$ we have $f_{s}(x)=\tfrac{\alpha}{2}\,\|x\|_{2}^{2}-\beta\,s^{\top}x,$ whose unique minimizer on $\Omega_{s}$ solves $\alpha x-\beta s=0$, hence $x_{s}^{*}=(\beta/\alpha)\,s\in\Omega_{s}.$ Therefore all $2^{n}$ points $(\beta/\alpha)\,s$, with $s\in\{-1,1\}^{n}$, are global minimizers of the function $f$, and the minimum value is $\min_{x\in\mathbb{R}^{n}}f(x)=-\,{n\beta^{2}}/{(2\alpha)},$ which is independent of $s$.
+The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization, see [Libertietall2014] and [2] for a comprehensive survey on this subject.
 This construction is also connected to Example 4.4.
@@
 In particular, $f$ is convex on each ${\cal C}_{i}$, hence star–convex there, and thus piecewise star–convex with respect to the polyhedral partition $\{{\cal C}_{i}\}_{i=1}^{m}$.
-Next, observe that $f$ admits a DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
+Next, observe that $f$ admits a nonsmooth DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
 $\min_{1\leq i\leq m}\Bigl\{\tfrac{1}{2}\,x^{\top}Qx+b_{i}^{\top}x+c_{i}\Bigr\}=\tfrac{1}{2}\,x^{\top}Qx+\min_{1\leq i\leq m}\bigl\{b_{i}^{\top}x+c_{i}\bigr\}=\tfrac{1}{2}\,x^{\top}Qx-\max_{1\leq i\leq m}\bigl\{-b_{i}^{\top}x-c_{i}\bigr\}.$
-Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx)$ and $h(x):=\max_{1\leq i\leq m}\bigl{(}-b_{i}^{\top}x-c_{i}\bigr{)}$.
+Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx$ and $h(x):=\max_{1\leq i\leq m}\bigl(-b_{i}^{\top}x-c_{i}\bigr)$.
 Clearly $g$ is $C^{\infty}$ and convex, while $h$ is the pointwise maximum of affine functions, hence convex, polyhedral, and Lipschitz.
-A particularly relevant special case of (8) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $f(x)=\min_{1\leq i\leq m}\,\|x-a_{i}\|^{2},\qquad x\in\mathbb{R}^{n}.$
+A particularly relevant special case of (11) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $\varphi_{i}(x)=\tfrac{1}{2}(x-a_{i})^{\top}Q(x-a_{i})$.
+In this case, each cell $\mathcal{C}_{i}$ contains $a_{i}$, and $\min_{x\in\mathcal{C}_{i}}f(x)=\min_{x\in\mathcal{C}_{i}}\varphi_{i}(x)=\varphi_{i}(a_{i})=0,$ with the minimizer in $\mathcal{C}_{i}$} given by $x_{i}^{\ast}=a_{i}$.
+Therefore, all cells share the same minimum value 0.
 Next we show that a basic robustness property of piecewise–star–convexity is that the star–type inequality is preserved by adding a convex term.
 Proposition 4.7.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}.
-Then $f+g$ is piecewise–star–convex on $\cal C$} with respect to $\cal P$}.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}.
+Then $f+g$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$}.
 Proof.
@@
 $f\bigl(\lambda x_{x}^{*}+(1-\lambda)x\bigr)\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x),\quad\forall\,\lambda\in[0,1].$
-Since $g$ is convex on ${\cal C}_{i}$,
-$g\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x),\quad\forall\,\lambda\in[0,1].$
-Summing the inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$.
+Since $g$ is convex on ${\cal C}_{i}$, we have $g\bigl(\lambda x_{x}^{*}+(1-\lambda)x\bigr)\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x)$, for all $\lambda\in[0,1]$.
+Summing these inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$.
 As the argument holds for every $i\in I$ and $x\in{\cal C}_{i}$, the claim follows.
@@
 Remark 2.
-If your definition of piecewise–star–convexity explicitly requires the reference points to be minimizers of the function under consideration on each cell, then the statement above remains valid provided $(f+g)$ attains a (local) minimum on each ${\cal C}_{i}$ (e.g., when each ${\cal C}_{i}$ is compact or $(f+g)$ is coercive there).
-In that case one may choose the reeference points among $\arg\min_{{\cal C}_{i}}(f+g)$.
-The next example combines Example 3.5 with Proposition 3.8 to show that a widely used optimization objective is piecewise–star–convex.
-Example 3.9.
-Let $A\in\mathbb{R}^{n\times n}$ be a full rank column, equivalently, $Q:=A^{\top}A$ a positive definite matrix, let $b\in\mathbb{R}^{n}$, and let $\beta>0$.
-Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
-$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{\,2}-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$
-Expanding the quadratic term gives the equivalent representation
-$f(x)\;=\;\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}-c^{\top}x\;+\;\tfrac{1}{2}\|b\|_{2}^{2},\qquad Q:=A^{\top}A,\qquad c:=A^{\top}b.$ (9)
-Equivalently, write $f=h+g$ with $h(x):=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}$ and $g(x):=-\,c^{\top}x+\tfrac{1}{2}\|b\|_{2}^{2}$.
-By Example 3.5, $h$ is piecewise–star–convex on the orthant partition when $Q\succ 0$; since $g$ is convex (affine), Proposition 3.8 implies that $f=h+g$ is also piecewise–star–convex on the same partition.
+The Example 3.2 follows immediately from Example 4.5 and Proposition 4.7.
 Proposition 4.8.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is closed and convex, interiors are pairwise disjoint, and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is nonempty convex and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$.
 Let $A\in\mathbb{R}^{n\times n}$ be invertible and $b\in\mathbb{R}^{n}$.
-Define the affine map
+Define the affine map $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by
 $\Psi(z):=Az+b,\qquad z\in\mathbb{R}^{n},$
-the transformed domain and cells $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
-$\tilde{f}(z):=f(Az+b),\qquad z\in\tilde{\mathcal{C}}$.
+the transformed sets $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
+$\tilde{f}(z):=f(\Psi(z)),\qquad z\in\tilde{\mathcal{C}}$.
 Then $\tilde{f}$ is piecewise–star–convex on $\tilde{\mathcal{C}}$ with respect to the partition $\tilde{\mathcal{P}}=\{\tilde{\mathcal{C}}_{i}\}_{i\in I}$.
 Proof.
-(i) Since $A$ is invertible, $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$.
+Since the matrix $A$ is invertible, the function $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$.
 If $z\in\tilde{\mathcal{C}}$ then $x:=\Psi(z)=Az+b\in\mathcal{C}$, so $\tilde{f}(z)=f(x)$ is well defined.
-Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is closed and convex; interiors remain pairwise disjoint and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$.
+Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is convex and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$.
 Thus $\tilde{\mathcal{P}}$ is a partition of $\tilde{\mathcal{C}}$.
-(ii) For each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$. We claim
+In addition, for each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$.
+We claim
 $\tilde{X}_{i}^{*}=A^{-1}(X_{i}^{*}-b).$ (12)
@@
 Equivalently, $z=A^{-1}(x-b)$ with $x\in X_{i}^{*}$, proving (12).
-(iii) Fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$.
-Because $f$ is piecewise–star–convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
+Now, fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$.
+Because $f$ is piecewise-star-convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
 $f\big((1-\lambda)x+\lambda x_{x}^{*}\big)\leq(1-\lambda)f(x)+\lambda f(x_{x}^{*}).$ (13)
-Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (10), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$,
+Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (12), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$, we conclude that
 $\Psi\big((1-\lambda)z+\lambda z_{z}^{*}\big)=(1-\lambda)\Psi(z)+\lambda\Psi(z_{z}^{*})=(1-\lambda)x+\lambda x_{x}^{*}.$
@@
 Fix $i\in I$ and $x\in{\cal C}_{i}$.
-Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ (a minimizer on the cell ${\cal C}_{i}$; hence, a local minimizer for $f$ on $\cal C$}) such that
+Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ a minimizer on the cell ${\cal C}_{i}$ such that
 $f(x_{x}^{*})-f(x)\ \geq\ f^{\circ}(x;\,x_{x}^{*}-x)\ \geq\ v^{\top}(x_{x}^{*}-x)\qquad\forall\,v\in\partial_{c}f(x),$ (14)
@@
 $f\bigl(x+\lambda(x_{x}^{*}-x)\bigr)\ \leq\ (1-\lambda)f(x)+\lambda f(x_{x}^{*}),\qquad\forall\,\lambda\in[0,1].$
-Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus (e.g., Theorem 2.1), $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (12).
+Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus, e.g., Theorem 2.1, $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (14).
 ∎
@@
 Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$.
-Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary point of the constrained problem $\min_{x\in{\cal C}}f(x),$ that is,
-$v^{\top}(x-\bar{x})\ \geq\ 0\qquad\forall\,v\in\partial_{c}f(\bar{x}),\ \forall\,x\in{\cal C}.$ (13)
-If, in addition, $\bar{x}\in{\cal C}_{i_{0}}$ for some $i_{0}\in I$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{0}}$, namely, $\bar{x}\in X_{i_{0}}^{*}:=\arg\min_{x\in{\cal C}_{i_{0}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{0}}}f(x).$
+Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary for problem (1), i.e., there exists $v\in\partial_{c}f(\bar{x})$ such that
+$v^{\top}(x-\bar{x})\geq 0\qquad\forall\,x\in{\cal C}.$ (15)
+In addition, for any $i_{*}\in I$ such that $\bar{x}\in{\cal C}_{i_{*}}$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{*}}$, namely, $\bar{x}\in X_{i_{*}}^{*}:=\arg\min_{x\in{\cal C}_{i_{*}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{*}}}f(x).$
 Proof.
-Apply Proposition 3.11 with $x=\bar{x}$ and $i=i_{0}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ with
-$f(x_{\bar{x}}^{*})-f(\bar{x})\ \geq\ v^{\top}(x_{\bar{x}}^{*}-\bar{x})\qquad\forall\,v\in\partial_{c}f(\bar{x}).$
-By stationarity (13) (taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{0}}\subset{\cal C}$), the right-hand side is $\geq 0$, hence $f(x_{\bar{x}}^{*})\geq f(\bar{x})$.
-Since $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ minimizes $f$ on ${\cal C}_{i_{0}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$.
-Therefore $f(x_{\bar{x}}^{*})=f(\bar{x})$ and $\bar{x}\in X_{i_{0}}^{*}$.
+Apply Proposition 4.9 with $x=\bar{x}$, $v\in\partial_{c}f(\bar{x})$ and $i=i_{*}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{*}}^{*}$ with
+$f(x_{\bar{x}}^{*})-f(\bar{x})\geq v^{\top}(x_{\bar{x}}^{*}-\bar{x}).$
+By condition (15), taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{*}}\subset{\cal C}$, the right-hand side is nonnegative, hence we conclude that $f(x_{\bar{x}}^{*})\geq f(\bar{x})$.
+Since $x_{\bar{x}}^{*}\in X_{i_{*}}^{*}$ minimizes $f$ on ${\cal C}_{i_{*}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$.
+Therefore, we have $f(x_{\bar{x}}^{*})=f(\bar{x})$, which implies that $\bar{x}\in X_{i_{*}}^{*}$.
 ∎
 Remark 3.
-In the piecewise setting, $X_{i}^{*}$ consists of the global minimizers of $f$ restricted to the cell ${\cal C}_{i}$, which are in general only local minimizers on the whole set $\cal C$}.
-Corollary 3.12 refines the global statement for piecewise-star-convex functions by showing that Clarke stationarity implies cellwise optimality.
-Moreover, if the cell ${\cal C}_{i_{0}}$ containing $\bar{x}$ attains the global minimum value, i.e., $\min_{{\cal C}_{i_{0}}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}.
-This highlights the flexibility of piecewise-star-convexity: by allowing the domain to be partitioned into regions where the function behaves “locally convex,” one captures a broader class of DC objectives than with classical star–convexity, while still retaining descent inequalities sufficient to establish convergence of Frank–Wolfe variants.
-4 The DC optimization problem
-We are interested in solving the following constrained DC optimization problem
-$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (14)
-where ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact and convex set, $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function and $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex function possibly non-differentiable, with domain on $\mathbb{R}^{n}$.
-Throughout the paper we assume that the gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, i.e., there exists a Lipschitz constant $L>0$ such that
-  • (A)
-    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
-Since we are assuming that ${\cal C}\subset{\mathbb{R}}^{n}$ is a compact set, its diameter is a finite number defined by
-$\operatorname{diam}({\cal C}):=\max\left\{\|x-y\|:~x,y\in{\cal C}\right\}.$ (15)
-Since ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact, the study of problem (14) is bounded from below.
-Then, the optimum value of the problem (14) satisfies $\infty<f^{*}:=\inf_{x\in{{\cal C}}}f(x)$ and optimal set $\cal C$$^{*}$ is non-empty.
-According to Theorem 2.3, the first-order optimality condition for problem (14) is stated as
-$(\nabla g({\bar{x}})-{\bar{u}})^{T}(x-{\bar{x}})\geq 0,\qquad\forall{\bar{u}}\in\partial g({\bar{x}}),\quad\forall x\in{{\cal C}}.$ (16)
-In general, the condition (16) is necessary but not sufficient for optimality.
-A point ${\bar{x}}\in{{\cal C}}$ satisfying condition (16) is called a stationary point to problem (14).
-Consequently, all $x^{*}\in{{\cal C}}^{*}$ satisfies (16).
-We finish this section by presenting a variant of a classical example of DC optimization problem that fits the aforementioned requirements.
-Example 4.1.
-Let ${\cal C}_{i}\subset{\mathbb{R}}^{n}$ be an arbitrary set, for all $i=1,\ldots,m$.
-The square distance with respect to the set ${\cal C}_{i}$, denoted by $d^{2}_{{\cal C}_{i}}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is defined
-$d^{2}_{{\cal C}_{i}}(x):=\inf_{y\in{\cal C}_{i}}\|x-y\|^{2}.$ (17)
-In general, the function $d^{2}_{{\cal C}_{i}}$ is not convex. However, due to $\|x-y\|^{2}=\|x\|^{2}-2x^{T}y+\|y\|^{2}$, the distance function $d^{2}_{{\cal C}_{i}}$ becomes to
-$d^{2}_{{\cal C}_{i}}(x):=\|x\|^{2}-\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right),$ (18)
-which is a difference of two convex functions, see [23, Example 2.1.4].
-Indeed, $d^{2}_{{\cal C}_{i}}=g-h$, where $g(x)=\|x\|^{2}$ is a convex quadratic function and $h(x)=\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)$ is a convex function given by the supremum of affine functions $x\mapsto\ell_{y}(x):=2x^{T}y-\|y\|^{2}$, for $y\in{\cal C}_{i}$.
-Let us state the constrained generalized Fermat-Weber location problem with square distances.
-For that, take $\omega_{i}\geq 0$, $i=1,\ldots,m$, such that $\sum_{i=1}^{m}\omega_{i}=1$ and note that
-$\sum_{i=1}^{m}\omega_{i}d^{2}_{{\cal C}_{i}}(x)=\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right).$
-Thus, for a given constrained set compact and convex ${\cal C}\subset{\mathbb{R}}^{n}$, the DC version of the constrained generalized Fermat-Weber location problem is stated as follows
-$\begin{array}[]{c}\min_{x\in{{\cal C}}}\left(\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)\right).\end{array}$ (19)
-Finally, observe that the objective function in problem (19) satisfies all the assumptions of problem 14.
-Moreover, in certain cases the objective function of the generalized Fermat–Weber problem (19) is weak–star–convex.
-For example, if $m=n$ and ${\cal C}_{i}=\{e_{i},-e_{i}\}$, where $e_{i}$ denotes the $i$-th canonical vector of $\mathbb{R}^{n}$, then the objective becomes $f(x)=\|x\|^{2}-2\sum_{i=1}^{n}\omega_{i}|x_{i}|+1,$ which, by the same reasoning as in Example 3.5, can be shown to be weak–star–convex.
-5 Classical Frank–Wolfe algorithm with an adaptive stepsize
-In this section, we present the classic Frank–Wolfe algorithm with an adaptive stepsize to solve problem (14) and provide its convergence analysis.
-Although we assume condition (A) for $g$, the Lipschitz constant will not be used explicitly to determine the stepsize.
-Instead, the stepsize is computed adaptively using the scheme introduced in [3, 5], which estimates the Lipschitz constant on the fly.
-Our convergence analysis shows that, for piecewise-star-convexfunctions, possibly non–differentiable and non–convex, the method achieves an ${\cal O}(1/k)$ convergence rate for both the function values and the duality gap.
-It is worth noting that a variant of the classic Frank–Wolfe algorithm for solving problem (14) was proposed in [26], where the stepsize depends on a curvature/Lipschitz–type constant of the gradient of $f$.
-For non–convex objective functions, the convergence analysis in [28] shows that such a variant attains only an ${\cal O}(1/\sqrt{k})$ rate in terms of reaching a stationary point.
-Throughout this section, we assume access to a linear optimization oracle (LO oracle) that can minimize linear functions over the set $\cal C$}. The algorithm is stated below.
+In the piecewise setting, $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ consists of minimizers of $f$ restricted to the cell ${\cal C}_{i}$.
+Such points are global minimizers on ${\cal C}_{i}$, but they need not be minimizers of $f$ on the whole set $\cal C$} and may fail to be even local minimizers in $\cal C$}.
+This is precisely the situation in which piecewise star-convexity is useful, star-type inequalities are enforced within each cell with respect to $X_{i}^{*}$, even though they may break down across cell boundaries.
+Corollary 4.10 refines the global picture by showing that Clarke stationarity implies cellwise optimality, namely $\bar{x}\in X_{i}^{*}$ for the cell ${\cal C}_{i}$ containing $\bar{x}$.
+In particular, if that cell achieves the smallest cell minimum value, that is, if $\min_{{\cal C}_{i}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}.
+5 Frank–Wolfe algorithm
+In this section, we present the Frank–Wolfe algorithm for solving (1) under a Lipschitz-based adaptive stepsize rule.
+We also establish iteration-complexity guarantees under piecewise star-convexity, see [3, 5].
+Our analysis shows that, for piecewise star-convex functions with a nonsmooth DC structure, the method attains an ${\cal O}(1/k)$ rate for the function values (relative to a best cell minimum) and a corresponding ${\cal O}(1/k)$ bound for the Frank–Wolfe gap.
+For comparison, a related Frank–Wolfe variant for (1) was proposed in [23], with stepsizes based on a curvature/Lipschitz-type constant.
+For general nonconvex objectives, the analysis in [26] yields an ${\cal O}(1/\sqrt{k})$ rate in terms of stationarity.
+Throughout this section, we assume access to a linear optimization oracle (LO oracle) that minimizes linear functions over $\cal C$}. The algorithm is stated below.
 Algorithm 1.
@@
 Step 2.
-Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}(x_{k})$ as follows
-$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}(x_{k}):=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (20)
+Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}_{k}$ as follows
+$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{k}:=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (16)
 Step 3.
-If ${\omega}(x_{k})=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
-$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}(x_{k})|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (21)
+If ${\omega}_{k}=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
+$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{k}|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{k}|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (17)
 Step 4.
 If
-$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (22)
+$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}_{k}|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (18)
 then set $j_{k}=j$ and go to Step 5. Otherwise, set $j=j+1$ and go to Step 3.
@@
 Set $k\leftarrow k+1$, and go to Step 1.
-Let us describe the main features of Algorithm 1.
-First of all, note that due to $x^{0}\in{C}$, $p^{k}\in{C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and C{C} be convex set, it follows from inductive arguments and first equality in (23) that $(x^{k})_{k\in{\mathbb{N}}}\subset{C}$.
-Note that (20) implies that
-${\omega}(x_{k})\leq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall p\in{C}.$ (24)
-Consequently, if ${\omega}(x_{k})=0$, then $(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0$, for all $p\in{C}$, i.e., $x^{k}$ is a stationary point to problem (14).
-Note that letting $p=x^{k}$ in (24) we conclude that
-${\omega}(x_{k})\leq 0,\qquad k=0,1,\ldots.$ (25)
-In order to simplify the notations, from now on we will use the following two notations:
-${\omega}_{k}:={\omega}(x_{k}).$ (26)
-Since $\omega_{k}=0$ implies that $x^{k}$ is a stationary point to problem (14).
-Thus, in view of (25) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$.
-The next result shows that the backtracking line–search in Algorithm 1 is well defined, identifies the smallest $j_{k}$ in Step 4, and establishes a descent inequality for the sequence $(x^{k})_{k\in{\mathbb{N}}}$.
+The LO oracle provides $p^{k}\in{\cal C}$ and induces the direction $d^{k}:=p^{k}-x^{k}$, with $p^{k}:=p(x^{k},u^{k})$ as defined in in (5).
+The scalar quantity $\omega_{k}:=\omega(x^{k},u^{k})\leq 0$, where $\omega(x^{k},u^{k})$ as defined in (5) is the corresponding Frank–Wolfe gap and vanishes only at stationary points, see Proposition 3.1.
+Therefore, the algorithm stops whenever $\omega_{k}=0$.
+From now on, we assume
+$\omega_{k}<0,\qquad\qquad k=0,1,\ldots.$ (20)
+for all $k\geq 0$, so the method generates an infinite sequence $\{x^{k}\}_{k\in\mathbb{N}}\subset{\cal C}$.
+In fact, since the update of $x_{k+1}$ in  (19) is a convex combination of $x^{k}$ and $p^{k}$ with $\lambda_{k}\in(0,1]$, the convexity of $\cal C$} yields $x^{k}\in{\cal C}$ for every $k$.
+Next we apply Proposition 2.4 to show that the backtracking procedure in Step 2 and 3 terminates in finitely many trials at each iteration.
+Therefore, $j_{k}$, $\lambda_{k}$, and the update $L_{k+1}$ in (19) are well defined.
+The next proposition states this formally and records the key descent inequality used in the complexity analysis.
 Proposition 5.1.
@@
 $f(x^{k}+\lambda_{k}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}_{k}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}.$ (22)
-In addition, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. And the following inequality holds
+In addition, the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 is well defined. And the following inequality holds
 $f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k},\qquad\qquad k=0,1,\ldots.$ (23)
@@
 $f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega_{k}}|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L-{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}.$
-Hence, it follows from the last inequality that (22) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proofs the first statement.
+Hence, it follows from the last inequality that (18) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proves the first statement.
 Consequently, $j_{k}$ is well defined and is the smallest non-negative integer satisfying (21) and (22).
-Therefore, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined.
+Therefore, the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 is well defined.
 We proceed with the proof of (23), by showing the following inequality
 $|{\omega_{k}}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}\leq-\frac{1}{2}|{\omega_{k}}|\lambda_{k}.$ (26)
-For that we separately consider two cases: $\lambda_{k}=|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$ and ${j_{k}}=1$.
-In the former, we have $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}=|{\omega_{k}}|/{\lambda_{k}}$,which substituting into the left hand side of (32) yields an equality.
-If now $\lambda_{k}=1$, then it follows from (21) that $\lambda_{k}=1\leq|{\omega_{k}}|/({2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2})$ or equivalently ${2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|$ and $\lambda_{k}=1$, which again substituting into the left hand side of (32) yields the inequality.
+For that, we note that $0\leq\lambda_{k}\geq|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$, from which we obtain that $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|/{\lambda_{k}}$.
+Substituting this into the left hand side of (26) yields the inequality.
 Hence, (26) holds.
@@
 ∎
+5.1 Convergence analysis
+In this section we analyze Algorithm 1 under Assumption (A).
+We first show that the backtracking rule yields uniformly bounded Lipschitz estimates and a uniform lower bound on the accepted stepsizes.
+These bounds, combined with the sufficient decrease inequality in Proposition 5.1, imply monotone descent of $\{f(x^{k})\}_{k\in\mathbb{N}}$, $\lim_{k\to\infty}\omega_{k}=0$, and Clarke-stationarity of every accumulation point.
+If $f$ is piecewise-star-convex, any such limit point is a global minimizer on its cell and a local minimizer of $f$ whenever it lies in the interior of that cell.
 For simplifying the notations we define the following constants
 $\alpha:=2\,(L+L_{0})\,\operatorname{diam}({\cal C})^{2}>0.$ (27)
+The next lemma provides a uniform upper bound on the backtracking estimates produced, and yields a convenient lower bound on the accepted stepsizes $\lambda_{k}$.
 Lemma 5.2.
-The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
-$L_{0}\leq L_{k}\leq L+L_{0},\quad\qquad k=0,1,\ldots.$ (34)
-and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}}$ satisfies
-$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{\alpha}\right\},\quad\qquad k=0,1,\ldots.$ (35)
+Assume (A) and let $(L_{k})_{k\in\mathbb{N}}$ be generated by (19). Then
+$0<L_{k}\leq L+L_{0},\qquad k=0,1,\ldots.$ (28)
+Moreover, the stepsizes $(\lambda_{k})_{k\in\mathbb{N}}$ selected by Lipschitz-based adaptive stepsize satisfy
+$\lambda_{k}\ \geq\ \min\Bigl\{1,\ \frac{|{\omega}^{k}|}{\alpha}\Bigr\},\qquad k=0,1,\ldots.$ (29)
 Proof.
-We proceed by induction to prove (34).
-Since the Lipschitz constant $L>0$, (34) holds for $k=0$.
-Now, assume that (34) holds for some $k>0$.
-We separately consider two cases: $j_{k}=0$ and $j_{k}\geq 1$.
-First assume that $j_{k}=0$.
-In this case, Step 1 together with second equality in (23) imply that $L_{k+1}=2^{-1}L_{k}\geq L_{0}$.
-Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}=0$.
-Since we are under the assumption that (34) holds, we have $L_{k+1}=2^{-1}L_{k}<L_{k}\leq L+L_{0}$.
-Thus, we conclude that two inequalities in (34) holds for $k+1$ and $j_{k}=0$.
-Next, we assume that $j_{k}\geq 1$.
-Again, Step 1 together with second equality in (23) yield $L_{k+1}:=2^{j_{k}-1}L_{k}\geq L_{0}$.
-Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}\geq 1$.
-To prove the second one, we assume by contradiction that $L_{k+1}>L+L_{0}$.
-This condition and (23) imply that $2^{j_{k}-1}L_{k}>L$.
-Thus, the first part of Proposition 5.1 implies that $j=j_{k}-1$ satisfies (22).
-Since the condition $L_{k+1}>L+L_{0}$ together (23) also implies $2^{j_{k}-1}L_{k}>L_{0}$, we have a contradiction because Proposition 5.1 also implies that $j_{k}$ is the smallest non-negative integer satisfying (27) and (28).
-Therefore, (34) holds for $k+1$, which concludes the proof that (34) holds for all $k\in{\mathbb{N}}$.
-The inequality (35) is an immediate consequence of (15), (21), the second inequality in (34) and (33).
+We first prove (28).
+Since $L_{0}>0$, we have $0<L_{0}\leq L+L_{0}$.
+Assume that $0<L_{k}\leq L+L_{0}$ for some $k\geq 0$ and recall from (19) that $L_{k+1}=2^{\,j_{k}-1}L_{k}.$ If $j_{k}=0$, then $L_{k+1}=\tfrac{1}{2}L_{k}$, hence $0<L_{k+1}\leq L_{k}\leq L+L_{0}$.
+If $j_{k}\geq 1$, suppose by contradiction that $L_{k+1}>L+L_{0}$.
+Then $2^{\,j_{k}-1}L_{k}=L_{k+1}>L$, which implies that $2^{\,j_{k}-1}L_{k}\geq L$.
+By the first statement of Proposition 5.1, the backtracking inequality (18) holds for $j=j_{k}-1$.
+This contradicts the fact that $j_{k}$ is the first index accepted by the backtracking loop in Step 3.2 of Lipschitz-based adaptive stepsize procedure.
+Hence $L_{k+1}\leq L+L_{0}$, and in all cases $0<L_{k+1}\leq L+L_{0}$.
+By induction, (28) holds for all $k$.
+We now prove (29).
+By (17) with $j=j_{k}$,
+$\lambda_{k}=\min\Bigl\{1,\ \frac{|\omega_{k}|}{2^{\,j_{k}}L_{k}\,\|p^{k}-x^{k}\|^{2}}\Bigr\}.$
+Since $x^{k},p^{k}\in{\cal C}$, we have $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$ by (2).
+Moreover, from (19) we obtain $2^{\,j_{k}}L_{k}=2L_{k+1}$, and then (28) yields $2^{\,j_{k}}L_{k}=2L_{k+1}\leq 2(L+L_{0}).$ Combining these estimates gives
+$\lambda_{k}\geq\min\Bigl\{1,\ \frac{|\omega_{k}|}{2(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}\Bigr\}=\min\Bigl\{1,\ \frac{|\omega_{k}|}{\alpha}\Bigr\},$
+where $\alpha$ is defined in (27). This is (29) and the proof is concluded. ∎
+In the next theorem, we summarize the resulting convergence guarantees for Algorithm 1.
+Theorem 5.3.
+Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be generated by Algorithm 1. Then, the following statement holds:
+  • (i)
+    The objective values sequence $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing. Moreover, the Frank–Wolfe gap satisfies $\lim_{k\to\infty}\omega_{k}=0.$
+  • (ii)
+    Every accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$ is Clarke-stationary for (1);
+  • (iii)
+    If, in addition, $f$ is piecewise-star-convex on $\cal C$} with respect to a partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, then for any such limit point $\bar{x}$ and any $i_{*}\in I$ with $\bar{x}\in{\cal C}_{i_{*}}$ and $\bar{x}\in X_{i_{*}}^{*}$, i.e., $\bar{x}$ is a global minimizer of $f$ on its cell ${\cal C}_{i_{*}}$.
+Furthermore, if $\bar{x}\in{\rm int}\,{\cal C}_{i_{*}}$, where ${\rm int}\,{\cal C}_{i_{*}}$ denotes the interior of the set ${\cal C}_{i_{*}}$, then $\bar{x}$ is local minimizer of $f$.
+Proof.
+To prove item (i), first recall that $\omega_{k}\neq 0$ for all $k$, which implies that an infinite sequence is generated.
+By Proposition 5.1, the accepted stepsize satisfies the sufficient decrease inequality (23), and hence $0\leq\frac{1}{2}|{\omega}_{k}|\leq f(x^{k})-f(x^{k+1})$ for all $k=0,1,\ldots,$ which implies that $\{f(x^{k})\}$ is monotone nonincreasing.
+We proceed to prove that $\lim_{k\to\infty}\omega_{k}=0$.
+Let $\alpha>0$ be defined in (27).
+Combining the sufficient decrease (23) with the lower bound (29), we obtain
+$\tfrac{1}{2}\min\bigl\{|{\omega}_{k}|,\tfrac{1}{\alpha}|{\omega}_{k}|^{2}\bigr\}\leq f(x^{k})-f(x^{k+1}),\qquad k=0,1,\ldots.$ (30)
+Since $\{f(x^{k})\}_{k\in\mathbb{N}}$ is monotone nonincreasing and bounded below on $\cal C$}, it converges.
+Hence, we have $\lim_{k\to\infty}(f(x^{k})-f(x^{k+1}))=0$.
+Therefore, passing to the limit in inequality  (30) yields $\lim_{k\to+\infty}\min\bigl\{|{\omega}_{k}|,\tfrac{1}{\alpha}|{\omega}_{k}|^{2}\bigr\}=0,$ which implies that $\lim_{k\to\infty}\omega_{k}=0$.
+To prove item (ii), let $\bar{x}$ be a limit point of $\{x^{k}\}_{k\in\mathbb{N}}$ and take a subsequence $\{x^{k_{\ell}}\}_{\ell\in\mathbb{N}}$ such that $\lim_{\ell\to+\infty}x^{k_{\ell}}=\bar{x}$.
+Because $h$ is convex and $\cal C$} is compact, Proposition 2.7 ensures that $\bigcup_{x\in{\cal C}}\partial h(x)$ is bounded.
+Hence, $\{u^{k_{\ell}}\}_{\ell\in\mathbb{N}}$ is bounded and, passing to a further subsequence if needed, we may assume $\lim_{\ell\to+\infty}u^{k_{\ell}}=\bar{u}$, for some $\bar{u}\in\mathbb{R}^{n}$.
+By Proposition 2.6, the graph of subdifferential $\partial h$ is closed, which implies that $\bar{u}\in\partial h(\bar{x})$.
+Now fix any $x\in{\cal C}$.
+Since $p^{k_{\ell}}$ minimizes the linear functional in (16), we have
+$(\nabla g(x^{k_{\ell}})-u^{k_{\ell}})^{\top}(x-x^{k_{\ell}})\;\geq\;(\nabla g(x^{k_{\ell}})-u^{k_{\ell}})^{\top}(p^{k_{\ell}}-x^{k_{\ell}})=\omega_{k_{\ell}}.$
+Using that $\lim_{\ell\to+\infty}x^{k_{\ell}}=\bar{x}$, $\lim_{\ell\to+\infty}u^{k_{\ell}}=\bar{u}$, the continuity of $\nabla g$, and $\lim_{\ell\to+\infty}\omega_{k_{\ell}}=0$, we obtain that $(\nabla g(\bar{x})-\bar{u})^{\top}(x-\bar{x})\geq 0$, for all $x\in{\cal C}$, which is precisely the Clarke-stationarity condition for (1).
+Finally, we prove item (iii).
+Assume additionally that $f$ is piecewise-star-convex on $\cal C$} with respect to ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$.
+Let $\bar{x}$ be any accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$.
+Choose an index $i_{*}\in I$ such that $\bar{x}\in{\cal C}_{i_{*}}$, such an index exists because $\cal P$} covers $\cal C$}.
+Since item (ii) has been proved, $\bar{x}$ is Clarke-stationary for (1).
+Therefore, Corollary 4.10 applies and yields $\bar{x}\in X_{i_{*}}^{*}$, i.e., $\bar{x}$ is a global minimizer of $f$ over ${\cal C}_{i_{*}}$.
+It remains to show the local minimality statement when $\bar{x}\in{\rm int}\,{\cal C}_{i_{*}}$.
+Because ${\rm int}\,{\cal C}_{i_{*}}$ is open in $\mathbb{R}^{n}$, there exists $r>0$ such that $B(\bar{x},r)\subset{\cal C}_{i_{*}}$.
+For every $y\in{\cal C}\cap B(\bar{x},r)$ we have $y\in{\cal C}_{i_{*}}$, and since $\bar{x}$ minimizes $f$ on ${\cal C}_{i_{*}}$ it follows that $f(\bar{x})\leq f(y)$, for all $y\in{\cal C}\cap B(\bar{x},r)$.
+Hence $\bar{x}$ is a local minimizer of $f$ on $\cal C$}.
 ∎
-Next, we state iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1. Since the analysis closely follows that in [4], the proofs are omitted.
+We end this section by emphasizing that, compared with the schemes in [4, 23], the only algorithmic ingredient that differs is the stepsize selection performed by the backtracking procedure in Algorithm 1.
+In the next section, we use the resulting descent estimate to derive the corresponding convergence and iteration–complexity guarantees.
+5.2 Iteration-complexity analysis
+In this section, we derive iteration–complexity bounds for the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 when applied to the DC problem (1).
+We focus on objectives of the form f=ghf=g-h that are piecewise star-convex on $\cal C$}, where $g$ is convex and continuously differentiable with $L$-Lipschitz continuous gradient on $\cal C$}, and $h$ is convex and possibly nonsmooth.
+We begin with a structural lemma for finite convex partitions that requires only continuity of $f$ and monotonicity of the objective values.
+It states that, if the function values converge to the minimum attained on some cell of the partition, then after a finite number of iterations the sequence can only visit cells whose minimum value does not exceed that limit.
+This prefix–exclusion mechanism will be used later, once the descent analysis ensures that $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing.
+Lemma 5.4.
+Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let ${\cal P}=\{{\cal C}_{i}:i=1,\dots,m\}$ be a finite partition of $\cal C$} into nonempty convex sets.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be such that $f_{i}^{*}:=\min_{x\in{\cal C}_{i}}f(x)$ is well defined for every $i$.
+Let $\{x^{k}\}_{k\in\mathbb{N}}\subset{\cal C}$ be a sequence satisfying $f(x^{k+1})\leq f(x^{k})$ for all $k$, and suppose that, for some $i_{*}\in\{1,\dots,m\}$,
+$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{*}}^{*}.$ (31)
+Then there exists $N\in\mathbb{N}$ such that, for all $k\geq N$ and every index $i_{k}\in\{1,\dots,m\}$ with $x^{k}\in{\cal C}_{i_{k}}$,
+$f_{i_{k}}^{*}\;\leq\;f_{i_{*}}^{*}.$
+Proof.
+Define $\hat{I}:=\{\,i\in\{1,\dots,m\}:\ f_{i}^{*}>f_{i_{*}}^{*}\,\}.$ If $\hat{I}$$\emptyset$, then the conclusion holds by taking $N:=0$.
+Assume $\hat{I}\neq\emptyset$ and fix $i\in\hat{I}$.
+By (31), since $f_{i}^{*}>f_{i_{*}}^{*}$, there exists $N_{i}\in\mathbb{N}$ such that $f(x^{N_{i}})<f_{i}^{*}.$ Because the sequence $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing, it follows that
+$f(x^{k})\leq f(x^{N_{i}})<f_{i}^{*},\qquad\forall\,k\geq N_{i}.$
+If, for some $k\geq N_{i}$, we had $x^{k}\in{\cal C}_{i}$, then by definition of $f_{i}^{*}$ as the minimum of $f$ on ${\cal C}_{i}$ we would have $f(x^{k})\geq f_{i}^{*}$, contradicting the strict inequality above.
+Hence,
+$x^{k}\notin{\cal C}_{i},\qquad\forall\,k\geq N_{i}.$
+Now set $N:=\max_{i\in\hat{I}}N_{i}$, finite because $\hat{I}$ is finite.
+Then for every $k\geq N$ we have $x^{k}\notin{\cal C}_{i}$ for all $i\in\hat{I}$.
+Therefore, if $x^{k}\in{\cal C}_{i_{k}}$ with $k\geq N$, necessarily $i_{k}\notin\hat{I}$, which means $f_{i_{k}}^{*}\leq f_{i_{*}}^{*}$, and the desired inequality is proved.
+∎
+We now state a complexity result showing that, despite nonconvexity, piecewise–star–convex objectives achieve the same Frank–Wolfe iteration bound as in the convex setting.
 Theorem 5.5.
-$\lim_{k\to+\infty}{\omega}_{k}=0$.
-Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is a stationary point to problem (14).
-In addition, for every $N\in{\mathbb{N}}$, there holds
-$\min\left\{|{\omega}_{k}|:~k=0,1,\ldots,N\right\}\leq\max\left\{{{2(f(x^{0})-f^{*})}},\sqrt{2\alpha{(f(x^{0})-f^{*})}}\right\}\frac{1}{\sqrt{N+1}}.$
-Since $\cal C$}} is compact and the functions $g$ and $h$ are convex, we can define the following constants:
-$\Theta:=\min\left\{\frac{2}{{\bar{\Gamma}}\operatorname{diam}({\cal C})},\frac{1}{\alpha}\right\}\qquad{\bar{\Gamma}}:=\max_{x\in{{\cal C}}}\{\|\nabla g(x)\|+\|u\|:~u\in\partial h(x)\},\qquad\beta:=\frac{{\bar{\Gamma}}^{2}}{{\Theta L_{0}}}.$
-Since the proof of the next theorem follows a standard argument in the convergence analysis of line–search methods, we omit it.
-Theorem 5.4.
-Let $\epsilon>0$. Then, Algorithm 1 generates a point $x^{k}$ such that $|{\omega}_{k}|\leq\epsilon$, performing at most,
-$\left(\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}\right)\left(2+\log_{2}\left({\beta}\frac{1}{\epsilon^{2}}\right)\right)={\cal O}(|\log_{2}(\epsilon)|\epsilon^{-2})$
-evaluations of functions $f$, and $\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}+1={\cal O}(\epsilon^{-2})$ evaluations of gradients of $f$ and subgradients of $h$.
-5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
-In this section, we establish iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 when the objective function is a DC function f=ghf=g-h that is piecewise weak–star–convex, where the first component $g$ has a Lipschitz continuous gradient on ${\cal C}\subset{\mathbb{R}}^{n}$ and the second component $h$ is convex but not necessarily differentiable.
-As will become clear, the piecewise-star-convexity property plays a central role in our analysis.
-By combining this property with the descent estimates from Proposition 5.1, we obtain the following global complexity result for the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1.
-Theorem 5.5.
-Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a piecewise-star-convex on $\cal C$}, and let ${\cal P}=\{\,{\cal C}_{i}:i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets such that $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$.
-For each $i\in I$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
+Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let ${\cal P}=\{{\cal C}_{i}:i=1,\dots,m\}$ be a finite partition of $\cal C$} into nonempty convex sets.
+Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$}.
+For each $i\in\{1,\dots,m\}$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
 $f_{i}^{*}:=\min_{x\in{\cal C}_{i}}f(x)$
-be the corresponding minimum value. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exists $i_{0}\in I$ such that:
-$f(x^{k})-f_{i_{0}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k}$, for all $k=1,2,\ldots.$
-$\min_{\ell\in\left\{\lfloor\frac{k}{2}\rfloor+2,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-2}$, for all $k=3,4,\ldots,$ where $\lfloor k/2\rfloor=\max_{n\in\mathbb{N}}\left\{n\leq k/2\right\}.$
+be the corresponding minimum value. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exist $i_{*}\in\{1,\dots,m\}$ and $N\in{\mathbb{N}}$ such that:
+  • (i)
+    $f(x^{k})-f_{i_{*}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-N+1}$, for all $k\geq N$;
+  • (ii)
+    $\min_{\ell\in\left\{\lfloor{k-N+1}/{2}\rfloor+N+1,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-N-1}$, for all $k\geq N+2$.
 Proof.
-By (28) in Proposition 5.1, we have
-$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega(x^{k})|\,\lambda_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (36)
-On the other hand, by (21),
-$\lambda_{k}\;=\;\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big{\{}-|\omega(x^{k})|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big{\}}.$
-Hence, for any choice $\beta_{k}\in(0,1]$, the right-hand side of (36) at $\lambda=\lambda_{k}$ is bounded by its value at $\lambda=\beta_{k}$. Taking $\beta_{k}:=\tfrac{2}{k+2}$, we get
-$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
-Using $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$, together with (23) and Lemma 5.2 (which yields $\tfrac{2^{\,j_{k}}L_{k}}{2}\leq L+L_{0}$), we obtain
+Using (22) in Proposition 5.1 and recalling that $\omega_{k}\leq 0$, we obtain that
+$f\!\big(x^{k}+\lambda_{k}(p^{k}-x^{k})\big)\leq f(x^{k})-|\omega_{k}|\lambda_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (32)
+On the other hand, by definition of the Lipschitz-based adaptive stepsize $\lambda_{k}$, we have
+$\lambda_{k}=\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big\{-|\omega_{k}|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big\}.$
+Hence, for any $\beta_{k}\in(0,1]$, in particular for $\beta_{k}:=\tfrac{2}{k+2}$, we conclude from the last equality that
+$|\omega_{k}|\,\lambda_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}\leq-|\omega_{k}|\,\beta_{k}+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
+Thus, combining this previous inequality with the inequality in (32), we conclude that
+$f(x^{k+1})\leq f(x^{k})-|\omega_{k}|\,\beta_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
+Since $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$ and $\frac{2^{\,j_{k}}L_{k}}{2}=L_{k+1}\leq L+L_{0}$, by (19) and Lemma 5.2, we obtain
 $f(x^{k+1})\leq f(x^{k})-|\omega_{k}|\,\beta_{k}+(L+L_{0})\,\operatorname{diam}({\cal C})^{2}\,\beta_{k}^{2}.$ (33)
-Since $\cal C$} is compact, the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1 is bounded.
+Set $A:=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$.
+Since $\cal C$} is compact, $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded and has accumulation points.
+By Theorem 5.3(i), $\{f(x^{k})\}_{k\in\mathbb{N}}$ is monotone nonincreasing and thus convergent.
 Let $\bar{x}$ be a limit point, and let $(x^{k_{\ell}})_{\ell\in\mathbb{N}}$ be a subsequence such that $\lim_{\ell\to+\infty}x^{k_{\ell}}={\bar{x}}$.
-Because ${\cal P}=\{{\cal C}_{i}:i\in I\}$ is a partition of $\cal C$}, there exists $i_{0}\in I$ with $\bar{x}\in{\cal C}_{i_{0}}$.
-In addition, Theorem 5.3 implies that $\lim_{k\to+\infty}{\omega}_{k}=0$ and $\bar{x}$} is a stationary point to problem (14).
-Hence, by (3.12), $\bar{x}$ is a minimizer of $f$ on ${\cal C}_{i_{0}}$.
-Let $f_{i_{0}}^{*}$ denote the minimum value of $f$ on ${\cal C}_{i_{0}}$.
-Thus, $\lim_{\ell\to+\infty}f(x^{k_{\ell}})=f({\bar{x}})=f_{i_{0}}^{*}$.
-Moreover, (29) in Proposition 5.1 implies that $\big{(}f(x^{k})\big{)}_{k\in\mathbb{N}}$ is monotone nonincreasing.
-Therefore,
-$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{0}}^{*},\qquad\text{and}\qquad f_{i_{0}}^{*}\leq f(x^{k})\quad\forall\,k\in\mathbb{N}.$ (38)
-Since $f$ is piecewise weak–star–convex on $\cal C$}, Proposition 3.11 and Theorem 2.2 ensure that, for any $x^{\star}\in X_{i_{0}}^{*}$,
-$f_{i_{0}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$
-where $u^{k}\in\partial h(x^{k})$. Using (20) and (26), the definition of $\omega_{k}$ gives $0\geq f_{i_{0}}^{*}-f(x^{k})\geq\omega_{k},$ hence
-$0\leq f(x^{k})-f_{i_{0}}^{*}\leq|\omega_{k}|\qquad\forall\,k\in\mathbb{N}.$ (39)
-Now apply Lemma 2.8 to the sequences
-$a_{k}=f(x^{k})-f_{i_{0}}^{*}\leq b_{k}=|\omega_{k}|,\qquad\beta_{k}=\frac{2}{k+2},\qquad A=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$
-The recursion (37) together with (39) satisfies the hypotheses of Lemma 2.8, yielding
-$f(x^{k})-f_{i_{0}}^{*}\leq\frac{4(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k},\qquad\forall k\geq 1,$
-and
-$\min_{\ell\in\{\lfloor k/2\rfloor+2,\ldots,k\}}|\omega_{\ell}|\leq\frac{16(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k-2}.\qquad\forall k\geq 3.$
-These are precisely items (i)–(ii) of Theorem 5.5. ∎
-Theorem 5.5 shows that, under piecewise–star–convexity, the exact–gradient Frank–Wolfe method with our stepsize achieves the convex-case rate: $f(x_{k})-f^{\star}=\mathcal{O}(1/k)$ and the optimality certificate $|\omega_{k}|$ decays at the same order.
-In contrast, the next section focuses on the finite–difference variant.
-We do not pursue complexity bounds there; instead we establish asymptotic convergence under mild consistency assumptions on the difference radius and the adaptive stepsizes: the gradient surrogate converges to $\nabla g$ along the iterates, the scheme is well posed, and every accumulation point is first-order stationary (equivalently, the Frank–Wolfe gap and $|\omega_{k}|$ vanish).
-Thus the Frank–Wolfe method provides projection-free, asymptotic guarantees compatible with the piecewise–star–convex framework, even without claiming the convex-case $\mathcal{O}(1/k)$ rate.
-6 Frank–Wolfe algorithm with finite difference
-In this section we introduce a projection-free Frank–Wolfe variant that replaces $\nabla g$ by a finite-difference surrogate with an adaptive stepsize, while retaining the linear minimization oracle of Algorithm 1.
-The scheme is designed to the DC model f=ghf=g-h from earlier sections, $g$ smooth convex and $h$ convex (via subgradients), and fits the DC piecewise–star–convexity framework developed in previous sections.
-Our analysis targets asymptotic convergence of the finite-difference method: under mild consistency conditions on the probing radii and stepsizes, the surrogate tracks $\nabla g$, the algorithm is well posed, and every accumulation point is first-order stationary (the Frank–Wolfe gap and $|\omega_{k}|$ vanish).
-We do not claim rates for this variant; Theorem 5.5 remains the $\mathcal{O}(1/k)$ benchmark for the exact-gradient case.
-Since piecewise–star–convexity objectives are convex within cells, stabilization in a cell yields local convex geometry, which motivates our stepsize choice.
-Let $\{e_{i}\}_{i=1}^{n}$ denote the canonical basis of $\mathbb{R}^{n}$.
-For a differentiable $g:\mathbb{R}^{n}\to\mathbb{R}$ and a stepsize $s>0$, define the (componentwise) forward finite-difference approximation of $\nabla g(x)$ by
-$D(x,s):=\bigl{(}D_{1}(x,s),\ldots,D_{n}(x,s)\bigr{)},\qquad D_{i}(x,s):=\frac{g(x+se_{i})-g(x)}{s},\quad i=1,\ldots,n.$ (40)
-The next lemma (see [38, (8.4), p. 195]) provides an upper bound on the error of approximating $\nabla g(x)$ by $D(x,s)$.
-Lemma 6.1.
-Assume that $g$ satisfies the assumption (A). Then, there holds
-$\|\nabla g(x)-D(x,s)\|\leq\frac{{\sqrt{n}L}}{2}s,\qquad\forall x\in\mathbb{R}^{n},\quad\forall s>0.$
-Proof.
-Setting $\nabla g(x)=(\partial_{1}g(x),\ldots,\partial_{n}g(x))$ and using (40) we have
-$\|\nabla g(x)-D(x,s)\|^{2}\leq\sum_{i=1}^{n}\left(\partial_{i}g(x)-D_{i}(x,s)\right)^{2}.$ (41)
-Applying Proposition 2.4 with $v=e_{i}$ and $\lambda=s$, we conclude that $D_{i}(x,s)-\partial_{i}g(x)\leq(Ls)/2$.
-Thus, (41) implies that $\|\nabla g(x)-D(x,s)\|^{2}\leq n((Ls)/2)^{2}$, which implies the desired inequality.
-∎
-The statement of the Frank–Wolfe algorithm to solve problem (14) with finite difference approximating the gradient $\nabla g$ is as follows:
-Algorithm 2.
-Frank–WolfeC,f:=g-h algorithm with finite difference
-Step 0.
-Select $x^{0},x^{1}\in{{\cal C}}$, $x^{0}\neq x^{1}$ and $L_{1}>0$. Set $k=1$.
-Step 1.
-Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $i=0$.
-Step 2.
-Set
-$s_{ij}=\frac{2L_{1}}{2^{i+j}L_{k}\sqrt{n}}\|x^{k}-x^{k-1}\|.$ (42)
-Step 3.
-Use an “LO oracle” to compute an optimal solution $p^{ij}$ and the optimal value ${\omega}_{ij}(x_{k})$ as follows
-$p^{ij}\in{\rm argmin}_{p\in{{\cal C}}}(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{ij}(x_{k}):=(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{ij}-x^{k}).$ (43)
-Step 4.
-If ${\omega}_{ij}(x_{k})<0$, then set $i_{k}=i$ and go to Step 5. Otherwise, set $i\leftarrow i+1$, and go to Step 2.
-Step 5.
-Compute the step size $\lambda_{j}\in(0,1]$ as follows
-$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{i_{k}j}(x_{k})|}{{2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{i_{k}j}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda^{2}\right\}.$ (44)
-Step 6.
-If
-$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}-\frac{2^{j}L_{k}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$ (45)
-then set $j_{k}=j$ and go to Step 7. Otherwise, set $j\leftarrow j+1$ and go to Step 2.
-Step 7.
-Set $\lambda_{k}:=\lambda_{j_{k}}$, $p^{k}:=p^{i_{k}j_{k}}$, define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
-$x^{k+1}:=x^{k}+\lambda_{k}(p^{k}-x^{k}),\qquad\quad L_{k+1}:=2^{j_{k}-1}L_{k}.$ (46)
-Set $k\leftarrow k+1$, and go to Step 1.
-Let us describe the main features of Algorithm 2.
-First of all, note that due to $x^{0},x^{1}\in{\cal C}$, $p^{k}\in{\cal C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and $\cal C$} be convex set, it follows from inductive arguments and first equality in (46) that $(x^{k})_{k\in{\mathbb{R}}}\subset{\cal C}$.
-Note that (43) implies that
-${\omega}_{ij}(x_{k})\leq D(x^{k},s_{ij})^{T}(p-x^{k}),\qquad p\in{\cal C}.$ (47)
-Hence, considering that $x^{k}\in{\cal C}$, letting $p=x^{k}$ in the last inequality we conclude that
-${\omega}_{ij}(x_{k})\leq 0.$ (48)
-In order to simplify the notations, from now on we will use the following two notations:
-$s_{k}:=s_{i_{k}j_{k}},\qquad\qquad{\omega}_{k}:={\omega}_{i_{k}j_{k}}(x_{k}).$ (49)
-Proposition 6.2.
-Assume that in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$. Then,
-$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0,\qquad\forall p\in{{\cal C}},$ (50)
-i.e., $x^{k}$ is a stationary point to problem (14).
-Proof.
-If in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$, then it follows from (47) that
-$0\leq(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall i\in{\mathbb{N}},\quad\forall p\in{{\cal C}}.$ (51)
-Since (42) implies that $\lim_{i\to+\infty}s_{ij}=0$, it follows from (40) that $\lim_{i\to+\infty}D(x^{k},s_{ij})=\nabla g(x^{k})$. Hence, taking limit in (51) we conclude that (50) holds. ∎
-In view of Proposition 6.2, (48) and (49) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$.
-6.1 Well-definedness
-Herein, we establish that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined.
-Lemma 6.3.
-Let $s_{ij}$ be define in (42) and assume that ${2^{j}L_{k}}\geq 2L$. Then, there holds
-$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{L_{1}}}{2}\|x^{k}-x^{k-1}\|.$
-Proof.
-By applying Lemma 6.1 with $x=x^{k}$ and $s=s_{ij}$ we obtain that
-$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{\sqrt{n}L}}{2}s_{ij}$
-Therefore, using (42) and considering that ${2^{j}L_{k}}\geq 2L$ the desired inequality follows. ∎
-Lemma 6.4.
-Let $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$.
-Then, (45) holds and $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying the following two conditions $2^{j}L_{k}\geq 2L_{1}$ and (45).
-Consequently, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined and satisfies the following inequality:
-$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$ (52)
-Proof.
-It follows from Proposition 2.4 that
-$g(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (53)
-On the other hand, considering that $u^{k}\in\partial h(x^{k})$ it follows from Proposition (2.5) that
-$h(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{i_{k}j}-x^{k}\rangle$ (54)
-Thus, combining (53) with (54) and taking into account the equality in (43) and (48), some algebraic manipulations shows that
-$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (55)
-To proceed with the proof we need first to prove that for $\lambda_{j}$ defined in (44) we have
-$|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}\leq-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}.$ (56)
-For that we separately consider two cases: $\lambda_{j}=|{\omega}_{i_{k}j}(x_{k})|/{(2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2})}$ and $\lambda_{j}=1$.
-In the former, we have $2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2}=|{\omega}_{i_{k}j}(x_{k})|/{\lambda_{j}}$,which substituting into the left hand side of (56) yields an equality.
-If now $\lambda_{j}=1$, then it follows from (44) that $\lambda_{j}=1\leq|{\omega}_{i_{k}j}(x_{k})|/({2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2})$ or equivalently ${2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}\leq|{\omega}_{i_{k}j}(x_{k})|$ and $\lambda_{j}=1$, which again substituting into the left hand side of (56) yields the inequality.
-Hence, (56) holds.
-Therefore, (55) becomes
-$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (57)
-Since ${2^{j}L_{k}}\geq 2(L+L_{1}/2)>L$, we can apply Lemma 6.3 to obtain that
-$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}.$
-Hence, considering that $2\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}\leq\|x^{k}-x^{k-1}\|^{2}+\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}$, we have
-$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
-Thus, combining the last inequality with (57) and considering that $\lambda_{j}<1$, we conclude that
-$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{L+L_{1}/2-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
-Considering that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$, the last inequality implies that (45) holds.
-Since the first trial $j$ in (45) is defined in Step 1, i.e., $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $2^{j+1}L_{k}\geq 2^{j}L_{k}$ we conclude that $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying $2^{j}L_{k}\geq 2L_{1}$ and (45).
-Consequently, $(x^{k})_{k\in{\mathbb{N}}}$ is well defined, which proof the first three statements of the lemma.
-Therefore, (52) follows from the definition of $j_{k}$ in Step 6, (45) and (46), which concludes the proof.
-∎
-Since the proof of the next lemma is similar to Lemma 5.2 it will be omitted.
-Lemma 6.5.
-The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
-$L_{1}\leq L_{k}\leq 2(L+L_{1}),\quad\qquad k=1,2,\ldots.$ (58)
-and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies
-$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}}\right\},\quad\qquad k=1,2,\ldots.$ (59)
-6.2 Convergence analysis
-In this section we analyze the convergence properties of the sequence $(x^{k})_{k\in{\mathbb{N}}}$.
-For that, we assume that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is infinite.
-Lemma 6.6.
-The sequence $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}}$ is monotone and non-increasing. Moreover, there holds:
-$\lim_{k\to+\infty}\|x^{k}-x^{k-1}\|=0.$ (60)
-As a consequence, $\lim_{k\to+\infty}f(x^{k})={\bar{f}}$ for some ${\bar{f}}\geq f^{*}$.
-Proof.
-Since $L_{k}\geq L_{1}$, for all $k\in{\mathbb{N}}$, it follows from Lemma 6.4 that
-$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$
-which, taking into account that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, yields
-$f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\leq f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots..$
-which implies the first statement of the lemma.
-We proceed to prove the second statement.
-For that we first note that Lemma 6.5 implies that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$.
-Thus, using again that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, it follows from (52) that
-$\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\leq f(x^{k})-f(x^{k+1}),\qquad k=1,2,\ldots.$ (61)
-Let $\ell\in{\mathbb{N}}$ be such that $\ell>1$. Summing up the inequality (61) from $1$ to $\ell$ we obtain
-$\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}\leq f(x^{1})-f^{*}.$ (62)
-Considering that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$, some algebraic manipulations show that
-$\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}+\frac{L_{1}}{4}\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}.$
-Combining (62) with the last inequality we obtain that
-$\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\frac{4}{L_{1}}\Big{(}f(x^{1})-f^{*}+\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}\Big{)}.$
-Since the last inequality holds for all $\ell\in{\mathbb{N}}$, we have $\sum_{k=1}^{+\infty}\|x^{k+1}-x^{k}\|^{2}\leq+\infty$.
-Thus, (60) holds.
-To prove the last statement, we first note that due to $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in\mathbb{N}}$ be a non-increasing sequence and bounded from below by $f^{*}$, it must be convergent.
-Thus, we set ${\bar{f}}=\lim_{k\to+\infty}\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)$.
-Hence, considering that
-$f(x^{k})=f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}-({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\geq f^{*},$
-by using (60) we conclude that $\lim_{k\to+\infty}f(x^{k})={\bar{f}}\geq f^{*}$ and the proof is concluded. ∎
-For simplifying the notations we define the following constants
-${\bar{\alpha}}:=2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}>0.$ (63)
-Lemma 6.7.
-The following inequality holds:
-$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},$ (64)
-for all $k=1,2,\ldots$.
-Proof.
-Combining inequalities (52) with (58) and (59), and also using (63), we conclude that
-$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\min\left\{1,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$
-for all $k=1,2,\ldots$. Thus, (64) is an immediate consequence of the previous inequality. ∎
-In the next theorem we will use the following constant
-$\Gamma:=2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}.$ (65)
-Theorem 6.8.
-The following statements hold:
-  • (i)
-    $\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is a stationary point to problem (14);
-  • (ii)
-    for every $N\in{\mathbb{N}}$, there holds
-    $\min\left\{|{\omega}_{k}|:~k=1,\ldots,N\right\}\leq\max\left\{{\Gamma},\sqrt{{\bar{\alpha}}\Gamma}\right\}\frac{1}{\sqrt{N}}.$
-Proof.
-To prove the item (i), we first note that (64) is equivalent to
-$\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2}\Big{)}-\Big{(}f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\Big{)},$ (66)
-for all $k=1,2,\ldots$.
-It follows from Lemma 6.6 that $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}^{*}}$ monotone and non-increasing.
-Thus, considering that it is also bounded from below by $f^{*}$, we conclude that it converges.
-Hence, taking limit in (66) we have
-$\lim_{k\to+\infty}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}=0,$
-which implies the first statement of item (i).
-For proving the second statement of item (i), take $\bar{x}$} a limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ and subsequence $(x^{k_{i}})_{i\in{\mathbb{N}}}$ such that $\lim_{i\to+\infty}x^{k_{i}}={\bar{x}}$.
-Since $\cal C$}} is a compact set the sequence $(u^{k})_{k\in{\mathbb{N}}^{*}}$ is bounded.
-Thus, without loss of generality we assume that $\lim_{i\to+\infty}u^{k_{i}}={\bar{u}}$.
-On the other hand, Lemma 6.6 together with (42) imply that
-$\lim_{k\to+\infty}s_{k}=0.$ (67)
-Let $(s_{k_{i}})_{i\in{\mathbb{N}}^{*}}$ subsequence of $(s_{k})_{k\in{\mathbb{N}}^{*}}$. Thus, we conclude from (40) that
-$\lim_{i\to+\infty}D(x^{k_{i}},s_{k_{i}})=\nabla g({\bar{x}}).$ (68)
-Finally, using (43) we conclude that
-$(D(x^{k_{i}},s_{k_{i}})-u^{k_{i}})^{T}(p-x^{k_{i}})\geq{\omega}_{k_{i}},\qquad\forall p\in{{\cal C}}.$
-Therefore, taking limite in the last inequality and using (67) together (68) we obtain that
-$(\nabla g({\bar{x}})-{\bar{u}})^{T}(p-{\bar{x}})\geq 0,\qquad{\bar{u}}\in\partial g({\bar{x}}),\qquad\forall~x\in{{\cal C}},$
-which implies that $\bar{x}$} is a stationary point to problem (14). To prove the item (ii), take $N\in{\mathbb{N}}^{*}$. Thus, (66) implies that
-$\sum_{k=1}^{N}\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{1})+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\Big{)}-\Big{(}f(x^{N+1})+\frac{L_{1}}{2}\|x^{N}-x^{N-1}\|^{2}\Big{)}.$
-Hence, we conclude that
-$\min\left\{\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}:~k=1,\ldots N\right\}\leq\frac{1}{N}\left(f(x^{1})-f^{*}+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\right).$
-It follows from the last inequality that there exists ${\bar{k}}\in\{1,\ldots N\}$ such that
-$|{\omega}_{\bar{k}}|\leq\frac{1}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right),\qquad|{\omega}_{\bar{k}}|\leq\sqrt{\frac{{\bar{\alpha}}}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right)},$
-which, by using (65), implies the item (ii). ∎
+By Theorem 5.3(iii), there exists an index $i_{*}\in\{1,\dots,m\}$ such that $\bar{x}\in{\cal C}_{i_{*}}$ and $\bar{x}\in X_{i_{*}}^{*}$.
+Hence, we have
+$\lim_{k\to\infty}f(x^{k})=\lim_{\ell\to\infty}f(x^{k_{\ell}})=f(\bar{x})=f_{i_{*}}^{*},\qquad\text{and}\qquad f_{i_{*}}^{*}\leq f(x^{k})\ \ \forall\,k\in\mathbb{N}.$
+Now applying Lemma 5.4 to the monotone sequence $\{f(x^{k})\}$ and the finite partition $\cal P$}.we conclude that there exists $N\in\mathbb{N}$ such that, for every $k\geq N$ and every index $i_{k}$ with $x^{k}\in{\cal C}_{i_{k}}$,
+$f_{i_{k}}^{*}\leq f_{i_{*}}^{*}.$ (34)
+Fix $k$ and choose $i_{k}$ with $x^{k}\in{\cal C}_{i_{k}}$.
+Select $x_{k}^{*}\in X_{i_{k}}^{*}$ such that $f(x_{k}^{*})=f_{i_{k}}^{*}$.
+Since $f$ is piecewise-star-convex on $\cal C$}, Proposition 4.9 and Theorem 2.2 yield
+$f_{i_{k}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}),$ (35)
+for some $u^{k}\in\partial h(x^{k})$. Combining (34) with (35), we obtain, for all $k\geq N$, that
+$f_{i_{*}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$ (36)
+Moreover, since $p^{k}$ solves the LO oracle (16), we conclude that for every $x\in{\cal C}$, there holds
+$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x-x^{k})\ \geq\ (\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k})=\omega_{k}.$
+Applying this with $x=x_{k}^{*}$ and using (36), we obtain $0\geq f_{i_{*}}^{*}-f(x^{k})\geq\omega_{k}$ for all $k\geq N$. Hence,
+$0\leq f(x^{k})-f_{i_{*}}^{*}\leq|\omega_{k}|\qquad\forall\,k\geq N.$ (37)
+For each integer ${m}\geq 1$, define
+$a_{m}:=f\!\big(x^{N+{m}-1}\big)-f_{i_{*}}^{*},\qquad b_{m}:=\big|\omega_{\,N+{m}-1}\big|,\qquad\beta_{m}:=\frac{2}{{m}+1}.$ (38)
+Applying (33) with $A:=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$, k=N+m1k=N+{m}-1 and $\beta_{k}=\beta_{m}$, we get for all ${m}\geq 1$,
+$a_{{m}+1}\leq a_{m}-b_{m}\,\beta_{m}+A\,\beta_{m}^{2}.$
+Together with (37), we have $0\leq a_{m}\leq b_{m}$ for all ${m}\geq 1$. Therefore the hypotheses of Lemma 2.8 apply, yielding
+$a_{m}\leq\frac{4A}{{m}},\qquad\forall m\geq 1,\qquad\qquad\min_{\tau\in\{\lfloor{m}/2\rfloor+2,\ldots,{m}\}}b_{\tau}\leq\frac{16A}{{m}-2}\qquad\forall m\geq 3.$
+Returning to k=N+m1k=N+{m}-1 and considering (38), we conclude that for every $k\geq N$, there holds
+$f(x^{k})-f_{i_{*}}^{*}=a_{k-N+1}\leq\frac{4A}{k-N+1},$
+and for every $k\geq N+2$, there holds
+$\min_{\ell\in\{\lfloor(k-N+1)/2\rfloor+N+1,\ldots,k\}}|\omega_{\ell}|=\min_{\tau\in\{\lfloor(k-N+1)/2\rfloor+2,\ldots,k-N+1\}}b_{\tau}\leq\frac{16A}{k-N-1}.$
+Since $A=(L+L_{0})\operatorname{diam}({\cal C})^{2}$, these are exactly the bounds in items (i)(ii). ∎
+The next remark rewrites the estimates of Theorem 5.5 in an explicit $\varepsilon$–complexity form.
+In particular, it provides concrete iteration thresholds ensuring $\varepsilon$–accuracy for the objective residual, with respect to the limiting cell minimum, and for the Frank–Wolfe optimality certificate.
+Remark 4.
+Under the assumptions of Theorem 5.5, there exist $i_{*}\in\{1,\ldots,m\}$ and $N\in\mathbb{N}$ such that the following $\varepsilon$-complexity estimates hold.
+Given $\varepsilon>0$, any index $k$ satisfying
+$k\ \geq\ N-1+\frac{4}{\varepsilon}(L+L_{0})\operatorname{diam}({\cal C})^{2}$
+guarantees $f(x^{k})-f_{i_{*}}^{*}\ \leq\ \varepsilon.$ On the other hand, given $\varepsilon>0$, any index $k$ satisfying the following condition
+$k\ \geq\ N+1+\frac{16}{\varepsilon}(L+L_{0})\operatorname{diam}({\cal C})^{2}$
+guarantees the existence of an iterate index
+$\ell\in\Bigl\{\bigl\lfloor\tfrac{k-N+1}{2}\bigr\rfloor+N+1,\ldots,k\Bigr\}\qquad\text{such that}\qquad|\omega_{\ell}|\ \leq\ \varepsilon.$
+Equivalently, after $k$ iterations one can extract, from the last $\mathcal{O}(k)$ iterates, an $\varepsilon$-certificate for the Frank–Wolfe gap.
+In practice, we do not attempt to estimate $N$.
+Instead, we use the computable Frank–Wolfe gap as the stopping criterion and terminate as soon as $|\omega_{k}|\leq\varepsilon$ or, more robustly, when $\min_{0\leq t\leq k}|\omega_{t}|\leq\varepsilon$.
+Theorem 5.5 then guarantees that such a threshold will be reached after a finite number of iterations and provides an $\mathcal{O}(1/\varepsilon)$ worst–case bound, up to the transient index shift encoded by $N$.
+To orient the reader, we record a simple situation in which the complexity bounds of Theorem 5.5 apply from the first iterate, that is, one may take $N=0$.
+Remark 5.
+If all cellwise minimum values coincide, i.e., $f_{i}^{*}=f_{j}^{*}$ for every $i,j$, then in Theorem 5.5 one may take $N=0$.
+Indeed, no cell has a strictly larger minimum value, hence the index set $\hat{I}=\{\,i:\ f_{i}^{*}>f_{i_{*}}^{*}\,\}$ introduced in Lemma 5.4 is empty.
+Therefore, the exclusion step is not needed, and the comparison argument leading to (36) holds for every $k$ without any initial index shift.
+Consequently, the complexity bounds in items (i)–(ii) apply from the first iterate.
+In particular, this situation covers the globally star–convex case on $\cal C$}, a single–cell partition, for which the unique cell is $\cal C$} itself and $f_{1}^{*}=f^{*}_{\cal C}$.
+Hence, Theorem 5.5 reduces to the standard convex–type $\mathcal{O}(1/k)$ guarantees for the objective residual $f(x^{k})-f^{*}_{\cal C}$ and for the gap certificate $|\omega_{k}|$.
+Theorem 5.5 shows that, under piecewise–star–convexity, the Frank–Wolfe method equipped with our backtracking stepsize achieves the convex-case rate, namely, $f(x^{k})-f^{\star}=\mathcal{O}(1/k)$, and the optimality certificate $|\omega_{k}|$ decays at the same order.
+Thus, the proposed projection-free scheme enjoys the standard sublinear complexity guarantees while remaining compatible with the piecewise–star–convex framework.
 6 Conclusions
-We analyzed Algorithm 1 and its finite–difference variant Algorithm 2 within the DC piecewise-star-convex framework.
-For piecewise-star-convex objectives, the exact–gradient method attains the convex–case rate $\mathcal{O}(1/k)$ for both the function suboptimality $f(x_{k})-f^{\star}$ and the optimality certificate (duality/Frank–Wolfe gap).
-For the finite–difference method, we established asymptotic convergence and an $\mathcal{O}(1/\sqrt{k})$ decay of the gap.
-A full iteration–complexity theory for the finite–difference variant under piecewise-star-convex remains open.
-Promising directions include: (i) matching $\mathcal{O}(1/k)$ rates via tighter control of finite–difference bias and adaptive radii, (ii) guarantees that account for possible transitions across piecewise-star-convex cells without stabilization assumptions, and (iii) extensions to stochastic oracles, away–steps, and line–search variants.
+We analyzed Algorithm 1 within the DC piecewise–star–convex framework.
+In particular, for functions that are piecewise–star–convex on a compact convex set, we showed that the proposed Frank–Wolfe scheme equipped with the backtracking stepsize attains the same sublinear iteration complexity as in the convex case, namely, the function-value error (measured against the appropriate cellwise minimum) and the Frank–Wolfe gap both decay on the order of $1/k$.
+These results extend convex-type complexity guarantees to a broad class of nonconvex and potentially nonsmooth DC objectives, complementing recent rates established for globally star–convex functions.
+Natural directions for further work include, identifying verifiable conditions under which the iterates enter (and remain in) a single cell so that the bounds hold from the first iteration, refining the analysis to better quantify the transient phase before the “active” cell is identified, and extending the guarantees to common Frank–Wolfe enhancements such as away–steps and pairwise steps, as well as other line-search mechanisms.
 Funding